### v2

[…]
tcb@breakable marginparsep has been altered.
topmargin has been altered.
marginparpush has been altered.
The page layout violates the ICML style. Please do not change the page layout, or include packages like geometry, savetrees, or fullpage, which change it for you. We’re not able to reliably undo arbitrary changes to the style. Please remove the offending package(s), or layout-changing commands and try again.
Training of Scaffolded Language Models
with Language Supervision: A Survey
Matthieu Lin 1,2 Jenny Sheng 2 Andrew Zhao 1
Shenzhi Wang 1 Yang Yue 1 Victor Shea-Jay Huang 3 Huan Liu 4 Jun Liu 4
[…]
${}^{1\,}$eap Lab, Tsinghua University ${}^{2\,}$, Tsinghua University ${}^{3\,}$hanghai AI Lab ${}^{4\,}$i’an Jiaotong University
linmatthieu@gmail.com, jsheng415@outlook.com, gaohuang@tsinghua.edu.cn, liuyongjin@tsinghua.edu.cn
footnotemark: footnotetext: ✉ Corresponding author(s)
This survey organizes the intricate literature on the design and optimization of emerging structures around post-trained LMs. We refer to this overarching structure as scaffolded LMs and focus on LMs that are integrated into multi-step processes with tools. We view scaffolded LMs as semi-parametric models wherein we train non-parametric variables, including the prompt, tools, and scaffold’s code. In particular, they interpret instructions, use tools, and receive feedback all in language. Recent works use an LM as an optimizer to interpret language supervision and update non-parametric variables according to intricate objectives. In this survey, we refer to this paradigm as training of scaffolded LMs with language supervision. A key feature of non-parametric training is the ability to learn from language. Parametric training excels in learning from demonstration (supervised learning), exploration (reinforcement learning), or observations (unsupervised learning), using well-defined loss functions. Language-based optimization enables rich, interpretable, and expressive objectives, while mitigating issues like catastrophic forgetting and supporting compatibility with closed-source models. Furthermore, agents are increasingly deployed as co-workers in real-world applications—such as Copilot in Office tools or software development. In these mixed-autonomy settings, where control and decision-making are shared between human and AI, users point out errors or suggest corrections. Accordingly, we discuss agents that continuously improve by learning from this real-time, language-based feedback and refer to this setting as streaming learning from language supervision.
1 Introduction
Figure 1: Non-parametric training of scaffolded LMs focuses on learning from language. Parametric training has excelled in learning from demonstration (supervised learning), exploration (reinforcement learning), and observation (unsupervised learning) by using well-defined loss functions. Instead, non-parametric training enables using a LM to optimize in the language space, allowing for efficient and interpretable learning with fuzzy objectives and rich textual feedback. As scaffolded LMs interact with digital systems operated by humans, mixed autonomy Cui et al. (2023) becomes increasingly important. We anticipate that learning from language will serve as a crucial framework for scaffolded LMs that learn across episodes Silver & Sutton (2025) (streaming learning) from rich user feedback in mixed autonomy. Importantly, it circumvents catastrophic forgetting Hadsell et al. (2020), which prevents parametric training from learning across episodes. We refer readers to Sec. 4.4 for further discussion.
Scaffolded language models embed language models (LMs) in a framework that extends their ability beyond traditional NLP tasks to perform open-ended digital automation tasks. Notably, they mark the beginning of a new type of programs for digital automation OpenAI (2018). It uses language to interface with humans, digital systems, and developer-defined code. This opens up the opportunity for mixed autonomy where humans and AI share control and decision-making responsibilities to achieve a common goal Wang et al. (2024c); Cui et al. (2023).
We define a scaffolded LM as a structure that integrates the LM into a multi-step process with tools. Specifically, the post-training interface of LM enables the developer to specify the system behavior in natural language Grattafiori et al. (2024). The developer specifies available tools, and the LM executes them by generating the function calls and input arguments. Tools Wang et al. (2024e) typically correspond to digital systems or developer-defined code. The scaffold Wu et al. (2023); Liu (2022); AI (2024); Khattab et al. (2023) refers to a surrounding structured code that manages control flow (multi-step process), maintains state (prompt to LM), and executes tools based on responses of the LM. Thus, the resulting scaffolded LM can be viewed as a semi-parametric model, where the LM is the parametric component. In particular, the scaffold enables an interpretable and efficient way to modify the behavior of the ‘program’. For instance, we can add new or modify tools corresponding to APIs to interact with the digital world, developer-defined code to perform computations, or even other entities such as the human or another (scaffolded) LM. As well as creating detailed prompts with intricate constraints for the LM. The design and optimization of scaffolded LMs requires significant trial and error by expert developers. This involves identifying the LM’s failure modes by running the scaffolded LM on examples of user queries and then addressing them one by one by refining the system prompt, input template, available tools, or the scaffold’s code itself.
Accordingly, recent works Yuksekgonul et al. (2025); Zhao et al. (2024a) have proposed to use an LM as an optimizer that interprets intricate objectives, execution traces, and rich execution feedback to generate updated non-parametric variables. In practice, scaffolded LMs are naturally exposed to a wide variety of rich language feedback, such as user feedback or execution errors, that reveal about the failure modes. Beyond learning from execution feedback, the LM optimizer also extracts insights or reusable tools from successful execution traces Wang et al. (2025c); Zhao et al. (2024a).
As shown in Fig. 1, training of non-parametric variables focuses on learning from language. Notably, parametric training excels in learning from demonstrations (supervised learning) Ouyang et al. (2022), exploration (reinforcement learning) DeepSeek-AI et al. (2025) , and observations (unsupervised learning)Brown et al. (2020), using well-defined loss functions LeCun et al. (2015). In contrast, optimization in the language space Cheng et al. (2023) provides a learning system with intricate objectives expressed in natural language that is efficient and interpretable. The training of non-parametric variables circumvents key limitations of gradient descent, such as catastrophic forgetting Hadsell et al. (2020), while being compatible with closed-source models Anthropic (2024b); OpenAI (2024). In this survey, we dub this framework training of scaffolded language model with language supervision.
This survey organizes the intricate literature of compound AI systems Zaharia et al. (2024), LM pipelines Khattab et al. (2023), prompt optimization Zhou et al. (2023), language agents OpenAI (2025); Yao (2024); Xi et al. (2023); Masterman et al. (2024), AI workflow optimization Cheng et al. (2024); Yuksekgonul et al. (2025); Yin (2024), and experiential learning agents Zhao et al. (2024a) under the concept of training of scaffolded language model with language supervision. It focuses on the multi-step aspect that involves optimization beyond the prompt of a single LLM call Zhou et al. (2023) to multiple LM calls and tools. Specifically, compound AI systems Zaharia et al. (2024); Lin et al. (2024) focus on the systemic perspective (the interconnection of components) and overlook the multi-step aspect. Meanwhile, the literature in language agents Xi et al. (2023) neglects the efficiency of LM pipelines (also called workflows) Khattab et al. (2023) that leverage the task structure to reduce the burden in decision-making and supervise the LM with clearer instructions.
To help readers navigate through the paper, we briefly introduce the organization of this survey. As shown in Fig. 2, The survey contains three sections.
  • Sec. 2 focuses on how the post-training interface of LMs interacts with the scaffold. We identify two types of scaffold: agent and workflow. And focus on important design aspect of agent and workflow. Finally, we reflect on the trend of recent benchmarks.
  • Then, Sec. 3 organizes the intricate literature that uses a LM as an optimizer to train the non parametric variables. We identify a trend whereby these methods focus on training more complex scaffold (where complexity is defined by how much heavy lifting is done by the scaffold v.s. the LM to solve the task).
  • Finally, our discussion in Sec. 4 points out a key capability missing from current scaffolded LM: the ability to continuously learn across episodes Silver & Sutton (2025); Yao (2025). This motivates a novel setting for training scaffolded LMs with language supervision. This discussion introduces recent progress in RL for LMs as well as post-training. It grounds our understanding of non-parametric training into the broader field. It highlights how current applications of non-parametric training focuses on an inappropriate setting, where RL from verifiable reward Lambert et al. (2025) is a much better alternative.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=10pt, l sep=10pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center , where level=1text width=14em,font=,, where level=2text width=14em,font=,, where level=3text width=10em,font=,, [ Training of Scaffolded Language Models With Language Supervision, ver, color=mypink!75!black, fill=mypink!100, text=black [ Scaffolded Language Models §2, color=myblue!75!black, fill=myblue!100, text=black [ Post-Trained
Language Models §2.1, color=myblue!75!black, fill=myblue!60, text=black ] [ Tools §2.2, color=myblue!75!black, fill=myblue!60, text=black ] [ Multi-step Process §2.3, color=myblue!75!black, fill=myblue!60, text=black [ Agent §2.3.1, color=myblue!75!black, fill=myblue!60, text=black ] [ Workflow §2.3.2, color=myblue!75!black, fill=myblue!60, text=black ] ] [ Benchmarks §2.4, color=myblue!75!black, fill=myblue!60, text=black ] ] [ Training with Language Supervision §3, color=myyellow!75!black, fill=myyellow!100, text=black [ Prompt Optimization §3.1, color=myyellow!75!black, fill=myyellow!60, text=black ] [ Experiential Learning §3.2, color=myyellow!75!black, fill=myyellow!60, text=black [ Insights §3.2.1, color=myyellow!75!black, fill=myyellow!60, text=black ] [ Tools §3.2.2,color=myyellow!75!black, fill=myyellow!60, text=black ] ] [ AutoDiff frameworks §3.3, color=myyellow!75!black, fill=myyellow!60, text=black [ Graph Abstraction §3.3.1, color=myyellow!75!black, fill=myyellow!60, text=black ] [ Graph Execution and
Optimization §3.3.2,color=myyellow!75!black, fill=myyellow!60, text=black ] ] ] [ Discussion §4, color=mygreen!75!black, fill=mygreen!100, text=black [ Post-Training §4.1, color=mygreen!75!black, fill=mygreen!60, text=black ] [ Learned LM Optimizer §4.2, color=mygreen!75!black, fill=mygreen!60, text=black ] [ Multi-Turn RL §4.3, color=mygreen!75!black, fill=mygreen!60, text=black ] [ Streaming Learning
with Language Supervision§4.4, color=mygreen!75!black, fill=mygreen!60, text=black ] ] ]
Figure 2: Organization of this survey. This survey introduces a unifying paradigm centered on learning from language using the semi-parametric capabilities of scaffolded language models. Despite the complexity and fragmentation of the literature across domains such as language agents, prompt optimization, experiential learning, and compound AI systems, we adopt a deliberately simple and clear organizational structure. This simplicity is not a limitation, but a reflection of our effort to distill a coherent framework from a rapidly evolving field. It enables us to outline key research opportunities in learning from language supervision.
2 Scaffolded Language Model
A scaffolded LM refers to an LM embedded in a multi-step process with tools. Such a structure allows for extending an LM’s utility to interact with digital systems in similar ways as humans do Wang et al. (2024c). While prior works have considered a meta-generation algorithm as an agent and employed an LM optimizer to generate its design Hu et al. (2024); Zhang et al. (2025); Saad-Falcon et al. (2024), our definition of a scaffolded LM does not include meta-generation techniques Welleck et al. (2024) such as best-of-n sampling or tree of thought as a scaffold. Meta-generation refers to search algorithms Silver et al. (2016) for reasoning tasks Cobbe et al. (2021) where each step consists of a call to an LM. Importantly, in scaffolded LMs, each call to a LM expects a tool response.
To provide readers with a practical understanding of how this structure allows LMs to perform sophisticated cognitive tasks and improve from non-parametric training, we organize this section as follows. We introduce the post-training interface of language models that allows interaction with tools and defines their behavior in natural language. Next, we define what we refer to as a tool. Furthermore, we show how the multi-step process can be controlled to implement priors into how the task is solved. Finally, we discuss benchmarks to understand the shift induced by scaffolded LMs. Specifically, we are excited by novel applications where humans share control and decision-making responsibilities with AI Windsurf .
2.1 Post-trained Language Models
MODEL INPUT
[…]
Figure 3: Example of prompt Grattafiori et al. (2024). We omitted the end-of-turn token. The prompt always begins with a system and a user message. As shown, the system prompt includes the available tools as a JSON schema. Additionally, we also specify the function calling format for using the tools. Each call to an LM requires appending an assistant header for completion (not shown here). After execution of the tools, we append the tool response, and the assistant continues with the next completion.
From a probabilistic perspective, given an input $\mathbf{x}=(x_{1},\dots,x_{T})$ (also called a prompt) consisting of a sequence of tokens $x_{t}\in\mathcal{V}$, a post-trained LM defines a distribution over output sequences $\mathbf{y}=(y_{1},\dots,y_{N})$ conditioned on $\mathbf{x}$. The LM assistant generates $\mathbf{y}$ in an autoregressive manner:
$p_{\theta}(\mathbf{y}\mid\mathbf{x})=\prod_{t=1}^{N}p_{\theta}(y_{t}\mid\mathbf{x},y_{<t}).$
Post-training defines a broad objective where the LM is trained to be helpful, harmless, and honest Grattafiori et al. (2024). Typically, we use a mixture of supervised finetuning and preference learning Grattafiori et al. (2024). Notably, the prompt $\mathbf{x}$ consists of tasks such as function calling and instruction following.
As shown in Fig. 3, the input $\mathbf{x}$ corresponds to a chat where the sources of each message (typically, system, user, assistant, and tool) are specified by special headers. At inference, a user provides a query $\mathbf{q}$ and we parse it into a chat $\mathbf{x}$ with a system message and the query, optionally with a pre-appended user input template. In principle, at each step, the chat is updated with new messages from the assistant or tool. The system message contains, for instance, the general instructions, safety guidelines, constraints, and tools available. Additionally, post-trained LMs use function calling to interact with tools in principled ways. In formula, we denote a scaffolded LM as $\pi_{\phi}(\mathbf{q})$, where $\phi$ refers to non-parametric variables, including the system prompt and the user input template for parsing $\mathbf{q}$ into $\mathbf{x}$.
2.2 Tools
We follow the definition of Wang et al. Wang et al. (2024e) and define an LM-used tool as a function interface to a computer program that runs externally to the LM, where the LM generates the function calls and input arguments in order to use the tool. In that sense, a tool can be virtually anything, including the interface to a digital system, a scratch pad for thinking Anthropic (2025), a request to the user LangChain (2025b), another LM OpenAI (2025), or developer-defined code. Notably, the function calling corresponds to a structured output from the post-trained LM that generates function calls and input arguments. Function calling provides a principled way to use tools in the sense that it allows the output of the LM to be parsed and provides a unified way to describe available tools to the LM. Thus, we follow previous work and use the terms function calling and tool interchangeably Grattafiori et al. (2024). It is important to note that the order of arguments in tools, as well as the order of tools when calling multiple tools, matter because post-trained LMs are autoregressive models McCoy et al. (2023); Grosse et al. (2023).
In practice, a post-trained LM comes with built-in tools or developer-defined code as tools. As we will see in Sec. 3, training of tools refers to updating developer-defined code. Built-in tools OpenAI (2025); Anthropic (2025) typically includes:
  • a web search API Nakano et al. (2022) allows the LM to tackle tasks with public data. The LM typically generates a response by citing the retrieved content.
  • a vector database Lewis et al. (2020); Izacard & Grave (2021); Ram et al. (2023); Anthropic (2024); Jiang et al. (2020) for storing data as embeddings that can be retrieved through top-$k$ retrieval to a query. It is often used for accessing private data or as a memory module Park et al. (2023).
  • a computer use tool OpenAI (2025) that interacts with a headless or headful browser to perform browsing tasks.
  • a thinking tool Anthropic (2025), unlike ReAct Yao et al. (2023b), defines reasoning itself as a callable tool, enabling structured, mid-process reflection that offers a more adaptable alternative—allowing the language model to interleave reasoning steps with multiple tool calls to maintain coherence and improve complex task handling.
2.3 Multi-Step Process
Figure 4: Multi-step process of a scaffolded LM. The distinctive trait between agent and workflow is that for each workflow step, the developer controls the system prompt and user instruction to pre-define the sub-task at that step. This makes workflows much more efficient and predictable than agents. However, workflows are engineered for specific scenarios. In this figure, we assume the case where the tool execution happens on the developer side. For built-in tools, tool execution happens on the model (API) side OpenAI (2025).
A scaffolded LM involves a multi-step process with tools, where at each step, the LM outputs tool calls. As shown in Fig. 4, guiding the model at each step with tailored prompts reduces its decision-making burden. Thus, in practice, a workflow involves pre-defining the number of steps and crafting appropriate user or system prompts for each, while an agent focuses on engineering the system prompt and the set of tools.
These approaches mainly differ in how much autonomy the LM has, and we can also refer to this autonomy as the degree of agenticness LangChain (2025a). An agent works well in unstructured scenarios, i.e., in cases where it is hard to anticipate all the scenarios that the LM might face. In contrast, a workflow approach works well when we know the solution path to a problem and can thus directly define each step required to solve it. A good example is solving GitHub issues in SWE-Bench Jimenez et al. (2024). A workflow such as Agentless Xia et al. (2024) solves GitHub issues with the following steps. Roughly, it first calls the LLM with the GitHub repository and issue description to provide a list of files that potentially contain the bug. Those files are then used to prompt the LLM to choose relevant content inside those files. Then, it asks the LLM to generate multiple patches. In comparison, an agent such as Openhands Wang et al. (2024c) provides the LM with a terminal and a code interpreter to interact with the repository.
2.3.1 Agents
Agents solve tasks by deciding at each step which tool to use and how to use it. They also determine when a task is solved based on their own interpretation of results and goals. An agent starts with a system message and a user message and generates the first response. Then, at each step, it appends to the input $\mathbf{x}$ the LM response and tool execution response, and generates the next response provided with the tool execution response. It continues until it decides to stop or it reaches the maximum number of steps, and (optionally) returns a response to the user. Note that while Fig. 3 appends the LM response and tool response under the special headers, we can also append them to the user message. Besides providing the right tool to our agent, an important design of the agent is how it generates function calls and input arguments. In particular, it affects the horizon as well as the task complexity. Traditionally, people have used structured formats such as JSON or XML to generate tool calls. Recently, CodeAct Wang et al. (2024b) proposes to use executable Python code as a unified way to call tools. As shown in Fig. 5, compared to structured formats like XML, code inherently supports both control flow (e.g., conditionals and loops) and data flow, allowing intermediate results to be stored and reused via variables when coordinating multiple tool calls. They Wang et al. (2024b) find that it achieves comparable or better performance on atomic tool use, while excelling in tasks requiring multiple calls to multiple tools in multi-turn interactions. However, it requires a sandboxed environment for code execution.
2.3.2 Workflows
Workflows Xia et al. (2024); Lu et al. (2024); Yamada et al. (2025) shine in applications where a solution path can be pre-determined. In contrast to agents, they are rendered more predictable and reliable by limiting the LM’s autonomy. At each step, the input $\mathbf{x}$ consists of a system message and a user message that is updated to specify the sub-task at each step. Thus, workflows consist of a pre-determined number of steps. It is more cost-efficient as it avoids any unnecessary exploration steps Xia et al. (2024). Unlike agents that scale with the number of steps Pan et al. (2024), workflows scale by generating multiple responses in parallel Snell et al. (2024); Welleck et al. (2024); Chen et al. (2024a). Specifically, by fixing the solution path, the output format at each step is predictable. Thus, workflows use meta-generation techniques Welleck et al. (2024) such as best of $n$ sampling or majority voting Wang et al. (2023b) to scale test-time compute. For instance, in Yamada et al. (2025), the AI scientist-v2 workflow can reach workshop-level papers by implementing tree-based methods on top of its previous version Lu et al. (2024). In particular, by generating multiple solutions at each step, it effectively generates multiple solution paths, and then uses heuristics (e.g., LLM as judge Yao et al. (2023a)) to select the best-performing path.
Figure 5: Generating function calls and input arguments. We show an agent that uses a tool with structured format or executable code Wang et al. (2024b). Provided with a system prompt to control how it interacts with the tool, the model either generates output that is easily parsed or directly generates Python code that executes the tool.
2.4 Benchmarks
As scaffolded language models (LMs) become increasingly integrated into real-world Windsurf ; Wang et al. (2024c); Cursor (2023); Microsoft (2025), there is growing momentum toward developing benchmarks that reflect mixed autonomy settings—scenarios where humans and AI share control and collaborate to achieve shared goals. In practice, these models are not isolated problem-solvers; they operate through the same APIs, UIs, and tools as human experts, and seldom succeed on the first try Rein et al. (2025); Wijk et al. (2024); Wei et al. (2025a); Yoran et al. (2024); Jimenez et al. (2024); Pan et al. (2024); Zhong et al. (2025); Jha et al. (2025); Starace et al. (2025). This necessitates multi-turn interactions, human-in-the-loop feedback, and iterative refinement—dynamics absent from traditional benchmarks focused on raw intelligence (e.g., math Olympiads or code generation).
To evaluate real-world effectiveness, recent benchmarks have shifted from measuring standalone task accuracy to assessing human-AI collaboration and productivity gains. For example, Tau-Bench Yao et al. (2024a) and Sweet-RL Zhou et al. (2025) simulate human feedback using LMs to better reflect the back-and-forth nature of co-working. Benchmarks like HCAST Rein et al. (2025) and RE-Bench Wijk et al. (2024) calibrate task difficulty against human effort—reporting metrics like “fraction of human pace” rather than raw accuracy. SWE-Bench Jimenez et al. (2024) and SWE-Lancer Miserendino et al. (2025) go further, grounding evaluation in software engineering tasks from GitHub or Upwork—tasks that are compensated in real-world settings and evaluated via functional test cases.
This benchmarking evolution reflects a shift: from measuring intelligence to measuring impact OpenAI (2018). The goal is no longer to assess whether a model solves a task perfectly in isolation, but how reliably and usefully it supports human workflows. Mixed autonomy systems must fluidly ratchet delegation up or down, adapt to under-specified tasks, navigate unclear success criteria, and coordinate across asynchronous streams of work. As highlighted by the METR report, such messiness is correlated with task length and reflects the realities of intellectual labor. Accordingly, we argue for new evaluation paradigms that go beyond single-turn metrics and instead quantify collaborative efficiency, reliability, and human-labor savings in dynamic, real-world environments.
3 Training with Language Supervision
We view scaffolded LM as semi-parametric models where the LM is the parametric component $\theta$ that is fixed. Instead, we consider the training of non-parametric variables $\phi$ that are represented in language or code. Thus, we denote our scaffolded LM as $\pi_{\phi}$, where we omit $\theta$ as it is fixed. As shown in Fig. 7, non-parametric variables $\phi$ include the system prompt, the user input template, developer-defined code as tools, or the scaffold’s code. As shown in Fig. 6, we use an LM as an optimizer to generate the updated variables $\phi$ according to an objective $J(\phi)$, an execution trace $\tau$, and optionally execution feedback $f$, all provided in natural language.
As in machine learning (ML), we assume a training, validation, and test dataset sampled from the same (unknown) distribution. During training, we iteratively sample data point(s) and update $\phi$. At inference, $\phi$ is fixed. Unlike in ML, where we train with Monte Carlo samples to prevent catastrophic forgetting Hadsell et al. (2020), an LM optimizer works with interpretable variables $\phi$ where the role of batching is still an open research question Yuksekgonul et al. (2025). The datasets $\mathcal{D}$ consist of natural language queries $\{\mathbf{q}\}$ corresponding to user input from which we obtain execution traces $\tau\sim\pi_{\phi}(\mathbf{q})$. During training, the execution traces come from any scaffolded LM $\pi_{\phi^{\prime}}$ or the current scaffolded LM $\pi_{\phi}$. In that sense, it shares similarity with RL where we have on/off-policy training, offline/online training Sutton & Barto (2018). However, the training dynamics significantly differ from those in RL. While RL is about finding good data (exploration) and imitating that good data (exploitation) Eysenbach et al. (2020), recent works in Sec. 3.3 learn only from bad execution traces. This is because failed execution traces often yield rich textual feedback, such as error messages or simulated user responses.
Figure 6: LM optimizers interpret fuzzy objectives and execution trace with rich execution feedback. We illustrate a simple example with an agent. In practice, the objective contains the variables to optimize, and the execution trace shows how these variables were used. Optionally, execution feedback can be provided.
Central to this paradigm is an optimizer implemented with an LM as opposed to bootstrapping successful execution traces as in-context learning examples Khattab et al. (2022); Von Oswald et al. (2023). We further discuss how this paradigm differs from other approaches that have employed an LM and non-parametric components for improving a scaffold LM. First, in Reflexion Shinn et al. (2023), the updates are not retrained across tasks and are only used for retrying a task. Second, CoALA Sumers et al. (2024) is an orthogonal approach that works at run-time with a specific scaffold. In particular, it requires changing the design of the agent with a memory module. This memory module not only serves as long-context but also for improving the agent over time. However, this requires the agent to make additional decisions on what and when to store or retrieve knowledge from this memory module. In training with language supervision, recent works focus on a setting where we are provided with a fixed dataset for training. Thus, it can be seen as amortizing the expensive inference of cognitive architecture Sumers et al. (2024) into a training procedure.
Our organization follows the structure outlined in Fig. 7. Tracing their development, these methods have evolved to optimize more complex execution traces. Automatic prompt optimization focuses on an LM (as opposed to a scaffolded LM). It covers various types of objectives for improving the prompt of a target LM. Experiential learning trains agents where the tools and system prompts matter. Accordingly, the execution traces consist of interactions between the LM and the tools, from which the LM optimizer extracts reusable tools or insights. Finally, AutoDiff frameworks focus on workflows, which resemble classical software pipelines that include LM calls at key stages. These frameworks focus on recording the execution trace as a graph. The LM optimizer then uses this graph for execution and optimization.
3.1 Prompt Optimization
In prompt optimization Li et al. (2025b), we update the instruction $\phi$ of an LM (as opposed to a scaffolded LM). We prompt an LM optimizer with an objective $J(\phi)$ and (optionally) a textual feedback $f$.
In its simplest form, we provide query and response examples and prompt the LM to generate the corresponding instruction Khattab et al. (2023); Zhou et al. (2023).
Provided with different initial instructions, other methods define the objective as generating variations. Each variation is evaluated on a validation set, and the top-performing versions are retained. In Zhou et al. (2023), the objective is to generate semantically similar instructions. Evolutionary methods Fernando et al. (2024); Guo et al. (2024) combine parts of multiple instructions and then introduce random changes; we call this process crossover and mutation. In OPRO Yang et al. (2024), we define the objective as generating instructions with a higher score using previous instructions, their scores, and a random sample of input-output pairs from the training data. To introduce fine-grained variations of instructions, SAMMO Schnabel & Neville (2024) uses a more granular representation of prompts by breaking them down into smaller sections, such as input formatting. This fine-grained structure allows SAMMO to adjust specific components, such as changing the input format from raw text to a structured format (e.g., XML). In a similar vain, ABO Ma et al. (2024) implements the instruction as step-by-step guidelines. Thus, the LM optimizer only updates the specific step that failed.
Figure 7: Evolution of LM optimization approaches. LM optimization approaches have evolved from prompt optimization, then agents, and finally workflows. In all approaches, an LM acts as an optimizer for generating updates on $\phi$. Top: prompt optimization focuses on LMs as opposed to scaffolded LMs. The LM optimizer uses a single or multiple LM calls. Middle Experiential Learning extends to agents and focuses on extracting insight from tasks and refining tools. The LM optimizer uses a single LM call to update them from successful/unsuccessful execution traces. Bottom: AutoDiff Frameworks focus on workflows where computation is represented as a graph. Using textual feedback, the LM optimizer can optimize nodes of the graph (e.g., system prompts).
Another line of work uses an LM to generate rich textual feedback $f$, and defines the objective as following the feedback $f$ Pryzant et al. (2023); Wang et al. (2024d); Chen et al. (2024c); Austin & Chartock (2024). Importantly, textual feedback is obtained from sample queries where the target LM fails. Additionally, these works also propose summarizing multiple pieces of feedback Chen et al. (2024c) or updates Austin & Chartock (2024).
Self-discover Zhou et al. (2024a) focuses on the output structure of the instruction. The objective is to generate a JSON format with keys representing distinct reasoning steps. By filling in each key in an autoregressive manner during generation, it guides the target LM’s reasoning steps.
3.2 Experiential Learning
We borrow the term experiential learning from ExpeL Zhao et al. (2024a) to refer to the training of agents with language supervision. The LM optimizer considers an objective $J(\phi)$ (e.g., “improve set of tools according to feedback”, “extract reusable tool from execution trace”), user queries $\mathbf{q}$, execution trace $\tau$, optionally a textual feedback $f$, and generates the updated variable $\phi^{\prime}$. In particular, the execution trace $\tau$ of agents consists of a sequence of tool calls/responses and may end with a response to the user.
As discussed in Sec. 2.3.1, tools play an important role in agents. As shown in Fig. 8, $\phi$ is either the set of tools or insights that are provided to the agent’s system prompt. Specifically, prior works leverage successful and unsuccessful execution traces with different objectives.
3.2.1 Insights Extraction
ExpeL Zhao et al. (2024a) starts by collecting a dataset for training consisting of user queries $\mathbf{q}$ and corresponding execution trace $\tau$. Using a Reflexion agent Shinn et al. (2023), multiple execution traces may be sampled for a query $\mathbf{q}$. In particular, the Reflexion agent allows the agent to retry the task upon failing, which allows the collection of unsuccessful and successful trajectories $(\tau^{\text{succ}},\tau^{\text{fail}})$ on the same task instance $\mathbf{q}$. The LM optimizer refines a list of insights from successful and unsuccessful execution traces $(\tau^{\text{succ}},\tau^{\text{fail}})$ for a given query $\mathbf{q}$ or from $k$ successful trajectories $(\tau_{1}^{\text{succ}},...,\tau_{k}^{\text{succ}})$. At inference, those insights are fixed and provided to the scaffolded LM through its system prompt.
Follow-up works on ExpeL Zhao et al. (2024a), such as AutoGuide Fu et al. (2024b), modify insights to be selectively added to the context at each step. Insights are specific to a state and contain a description of the state in which it is applicable. However, these methods require additional cost at inference since an LM summarizes the current state and then uses it to select the corresponding state-specific insight. AutoManual Chen et al. (2024b) modifies ExpeL Zhao et al. (2024a) to prevent distributional shift in offline RL, where at test-time the state distribution is different Wang et al. (2023a). However, their paper does not show strong evidence that training agents from language supervision suffers from distributional shift Wang et al. (2023a). Additionally, MSI Fu et al. (2024a) proposes to categorize each insight and retrieve the relevant one at test-time based on pre-defined rules.
ExpeL Zhao et al. (2024a) also explores new settings such as transfer learning, where we use insights learned from a source dataset $\{\mathbf{q}^{\text{source}}\}$ and transfer it to a target dataset $\{\mathbf{q}^{\text{target}}\}$. In parametric training Pouyanfar et al. (2018), it consists of initializing the weights with the model trained on the source dataset. Then, finetuning these weights on the target dataset. In ExpeL, they initialize with insights extracted from a source dataset. Then, the LM optimizer’s objective is to generate refined insights to the target task provided with an execution trace (from a human). Note that the update is done in a single shot thanks to the LM’s ability to interpret intricate objectives. Additionally, they manually inspect execution traces from the agent with and without the insights. For instance, in ALFWorld Shridhar et al. (2020), one insight would be “when searching for an item, consider its nature and its typical usage.” This leads the agent to update its belief Andreas (2022) and avoid unnecessary actions to find objects. Furthermore, they find that another insight, “if an attempt to interact with an item fails […] consider alternative actions or locations,” leads to self-correction. These findings are important, they showcase that simple high-level insights influence the behavior of agents in some predictable way.
Figure 8: Experiential learning focuses on tools and insights. Successful execution traces are used for extracting reusable tools Wang et al. (2025c) while unsuccessful execution traces are used to refine tools Zhang et al. (2024). Additionally, we can contrast successful and unsuccessful trajectories on the same query to extract insights Zhao et al. (2024a).
3.2.2 Tools Optimization
Tool optimization Zhang et al. (2024); Zhao et al. (2024c); Wang et al. (2025c) focuses on refining developer-defined code into more effective routines. The main challenge of tool optimization is ensuring that the skills are reusable and useful across episodes Wang et al. (2024a).
A related term to tool, but slightly more general is skill Wang et al. (2024f). It refers to both a textual representation of a sequence of action and developer-defined code as tool. The term “skills” emphasizes two aspects of tool optimization. First, we want to make tools more efficient, thereby achieving tasks with fewer steps. Second, similarly to unsupervised RL (URL) Zhao et al. (2022); Eysenbach et al. (2019), we want to extract the set of behaviors possible in the environment. Drawing an analogy with principal component analysis, before we are given any tasks, we want to know what is the basis of behavior that exists in some environment. Similarly, these methods Liu et al. (2025b); Wang et al. (2024f) stores skills into a vector database and retrieve corresponding skill at inference. For instance, in CER Liu et al. (2025b) skills are represented as a sequence of tool calls with a high-level description of the skill. In AWM Wang et al. (2024f), they explore different textual representation of skills. We can also understand these methods as ‘curating’ in-context samples for retrieval of similar experiences in ExpeL Zhao et al. (2024a). However, the textual representation represents a challenge for formal verification Wang et al. (2025c).
At each step, LearnAct Zhao et al. (2024c) and AgentOptimizer Zhang et al. (2024) sample an execution trace $\tau\sim\pi_{\phi}(\mathbf{q})$, and refine the set of tools for unsuccessful execution traces $\tau^{\text{fail}}$ LearnAct uses an execution feedback $f$ corresponding to execution errors. The LM optimizer’s objective is to address $f$ by either revising, adding a tool, or updating the system prompt with information about the tools. AgentOptimizer does not rely on the execution feedback $f$. The LM optimizer modifies the set of tools at each step by revising, removing, or adding a tool to solve the task that the model failed on. Moreover, they propose heuristics to account for the LM’s inability to correctly understand the failure modes and update tools. At each epoch, they evaluate the performance of the agent on the training data. If the performance is lower than the previous epoch, they roll back the set of tools. However, this is expensive and quickly leads to overfitting, where tools are not reusable.
In contrast, ASI Wang et al. (2025c) extracts reusable tools from a successful execution trace. One important challenge in extracting a reusable tool is to make sure that it is both useful and bug-free. AgentOptimizer Zhang et al. (2024) checks the performance on the training set with and without the updated tools by simulating again on the same query. An LM is used to rewrite the execution trace with only the extracted tools. Then, the agent attempts to finish the execution. If it successfully solves the task, then the extract tool is kept. In contrast to transfer learning for insights, ASI explores domain adaptation Ovadia et al. (2019) in web browsing. They find that the agent can effectively reuse tools and adapt tools to the new domain.
Similar to unsupervised RL Zhao et al. (2022), recent works propose to use an LM to both propose tasks and extract tools from rollouts on these tasks. In Voyager Wang et al. (2024a), the LM optimizer refines tools based on execution feedback and adds them to the skill set when successfully used to achieve a task. SkillWeaver Zheng et al. (2025) extracts tool directly from successful execution traces. Then, it produces test cases to check the long-term viability of the tool.
3.3 AutoDiff Frameworks
The term autodiff is inspired by deep learning frameworks Paszke et al. (2019) in the sense that it implements elementary operations for defining and optimizing workflows Yuksekgonul et al. (2025); Cheng et al. (2024); Zhou et al. (2024b); Yin (2024). The frameworks represent execution traces with a directed acyclic graph (DAG) where nodes correspond to variables or operations and edges represent dependencies. They roll out execution traces $\tau\sim\pi_{\phi}$ and obtain execution feedback $f$. Subsequently, prior works Yuksekgonul et al. (2025); Yin (2024); Zhou et al. (2024b); Wang et al. (2025a) use the chain rule to back-propagate $f$ through the graph. Each variable of the graph is updated by the LM optimizer according to its feedback. Alternatively, Trace Cheng et al. (2024) converts the DAG into text akin to a Python traceback. The LM optimizer updates variables in an autoregressive manner. We organize this section as follows. First, we cover graph abstraction to understand which variables of the workflow are optimized in different frameworks. Second, we focus on how execution traces are recorded and used for optimization.
3.3.1 Graph Abstraction
We refer to graph abstraction as the representation of nodes in the graph. This abstraction is foundational because it determines how workflows are formalized, traced, and subsequently optimized. There are two types of nodes in the graph: variables and operations. Variables correspond to non-parametric components of the workflow (e.g., system prompt) or intermediate values (result from operations). Operations correspond to computation done on variables, e.g., Python function, LM call. As an analogy with graph representation for neural networks, variables would be weights, and operations would be layers. However, different from neural networks, any node of the computational graph can be optimized, including operations Cheng et al. (2024). We distinguish between two lines of work, focusing on prompts or Python functions. It is important to note that in this survey, we focus on workflows within the scope of scaffolded LM. Thus, we do not consider applications of these frameworks for optimization instances (as opposed to workflows).
TextGrad Yuksekgonul et al. (2025) and other variants Yin (2024); Wang et al. (2025a); Zhou et al. (2024b) focus on textual variables. In particular, operations include concatenations and LM calls Yuksekgonul et al. (2025). Accordingly, trainable variables correspond to the system prompt or user input template of an LM call.
Trace Cheng et al. (2024) focuses on general-purpose workflows. Importantly, operations are not limited to LM calls, they can be arbitrary Python functions. In practice, we wrap Python functions with the Python decorator “bundle”. Additionally, they support the optimization of nodes corresponding to operations. In that sense, it goes beyond prompts to optimizing the workflow’s code.
Chain Rule
Formatting of execution trace
Figure 9: Graph execution and optimization using chain rule vs. formatting of execution trace. AutoDiff frameworks consist of three key steps: (1) forward computation, (2) backward feedback propagation, (3) variables update. Both approaches differ during their backward pass. Left: Using the chain rule, we compute a grad field on each variable corresponding to a feedback that addresses the feedback on its child variable. The LM optimizer updates each trainable variable following its grad field. Right: It performs a reverse topological sort on the computation graph starting from the output to format the prompt template of the optimizer. Together with the feedback, the LM optimizer generates trainable variables in an autoregressive manner.
3.3.2 Graph Execution and Optimization
User input Template $\phi$Input data $\mathbf{q}$concat($\phi$, $\mathbf{q}$)inputLM callresponseevaluate(output)feedback
Figure 10: Chain rule in TextGrad. Forward computation is represented with solid black arrows. Calling “feedback.backward()” (dashed red arrows, bent) applies the chain rule on the feedback $f$. It computes a gradient for each variable (in green), provided with gradient functions attached to operations (in yellow).
We refer to graph execution and optimization as the process of sampling an execution trace $\tau\sim\pi_{\phi}(\mathbf{q})$ and optimizing variables $\phi$ of the graph according to the execution feedback $f$. As shown in Fig. 9, it involves three key steps: (1) forward computation, (2) backward feedback propagation, (3) variables update. During forward computation, we collect an execution feedback $f$. Then, backward feedback propagation performs a reverse topological sort to either apply the chain rule on the execution feedback Yuksekgonul et al. (2025); Yin (2024); Wang et al. (2025a); Zhou et al. (2024b) or format a prompt template with the DAG Cheng et al. (2024). On the one hand, the chain rule enables structured credit assignment by propagating feedback through the graph in a fine-grained manner. On the other hand, encoding the entire execution trace as a single prompt template is more token-efficient.
In the context of optimization in the language space, the chain rule does not refer exactly to the differentiation of composite functions in calculus. Instead, it points to the idea of when a transformation is applied to the output of another transformation. Then, the overall effect on an input can be understood by systematically combining the influence of each transformation in sequence. As shown in Fig. 10 and following the graph abstraction presented in Sec. 3.3.1. Operations, including LM call and concatenation, create nodes and attach a grad function. Most importantly, the grad function contains a prompt for an LM that informs about that operation’s role. Similarly, each variable has a role. These roles are provided in text to account for heterogeneous variables and operations in workflows. Accordingly, the backward feedback propagation calls the attached grad function at each operation in the reverse order of the DAG. Each gradient is a textual feedback with respect to the child node of that operation. For each trainable variable, the LM optimizer’s objective follows the feedback as well as the variable’s role. Follow-up work extends Textgrad Yuksekgonul et al. (2025) in the following ways. AdalFlow Yin (2024) extends operations to retrievers Lewis et al. (2020). GASO Wang et al. (2025a) modifies the gradient operation to take into account neighborhood nodes.
A conceptually more efficient approach is Trace Cheng et al. (2024). At each step of the optimization process, the chain rule requires multiple calls to an LM. Furthermore, the LM optimizer also requires multiple calls to an LM. In contrast, Trace Cheng et al. (2024) formats a prompt template for the execution trace during backward feedback propagation. This means the LM optimizer generates all updated variables in an autoregressive manner, provided with the execution trace and feedback. Additionally, as discussed in Sec. 3.3.1, Trace Cheng et al. (2024) uses a different graph abstraction that allows for arbitrary Python functions as operations. As opposed to Textgrad Yuksekgonul et al. (2025), this allows the LM optimizer to also consider operations during training. In that sense, we can optimize the workflow’s code.
4 Discussion
In the context of rapid progress in the field of LMs, it is important to ground our discussion to the broader research in LMs, specifically on the recent breakthroughs in LMs focusing on training paradigms that elicit general capabilities Sutton (2019). In particular, with recent advancements in RL for LMs DeepSeek-AI (2025), the utility of non-parametric training can be questionable, and we bring forward this discussion. For instance, as base models become stronger, RL from verifiable rewards (RLVR) Lambert et al. (2025) has proven to be a much more attractive approach to prompt optimization based on RL Deng et al. (2022). To understand how the training of scaffolded LMs with language supervision fits into the broader literature of RL for LMs, we organize the discussion as follows:
  • We start our discussion with how post-training affects scaffolded LMs.
  • Then, we focus on how better post-trained models have enabled the development of LM optimizers trained with RL.
  • We notice how recent works in RLVR have overshadowed prompt optimization techniques for LMs, as they directly optimize the LMs and obtain strong generalization capabilities. This motivates us to discuss recent progress in multi-turn RL for scaffolded LMs.
  • Finally, we highlight a setting where non-parametric training with language supervision becomes particularly well-suited—namely, mixed autonomy, where scaffolded LMs collaborate with humans and receive real-time textual feedback. In these settings, interactions are not i.i.d., and tasks often span multiple episodes with recurring context. While parametric training (e.g., via gradient descent) struggles under such conditions due to issues like catastrophic forgetting Hadsell et al. (2020), non-parametric approaches—such as updating prompts, tools, or scaffold logic—offer a more adaptable alternative. By learning from language in context, these systems can avoid redundant errors, build on past successes, and adapt dynamically. As such, we argue that non-parametric training will play a central role in enabling scaffolded LMs to learn continuously across episodes and operate effectively in dynamic, real-world environments Silver & Sutton (2025).
4.1 Post-Training
Post-training plays a synergistic role with non-parametric training. Many capabilities in post-training aim at improving their steerability through non-parametric variables such as instruction following, long context, reasoning, and function calling. As post-training methods improve, LMs become more steerable OpenAI (2025a). In turn, this allows solving tasks not seen during post-training with intricate prompts. More importantly, in the context of scaffolded LMs as semi-parametric models, we think of post-training as encoding the default behavior of the scaffolded LM. This section presents the multitude ways in which the default behavior of the LM is shaped and interacts with non-parametric variables.
Character training is a post-training step aimed at enriching language models with personality traits beyond helpfulness, such as curiosity, humility, openness, and thoughtfulness Anthropic (2024a). Unlike safety-motivated RLHF, which focuses on the content that the model outputs, character training focuses on the manner with which models respond Lambert (2025). The overarching goal is to improve the quality of interaction, making models more enjoyable and trustworthy conversational partners Anthropic (2024a). However, the details surrounding character training for foundation models are quite elusive and under-documented. For Claude 3, character shaping is done with a variant of the Constitutional AI pipeline Bai et al. (2022). For Llama 3, persona is part of the consideration when collecting steerability examples, which is then used for RLHF Grattafiori et al. (2024).
It is plausible that character training for the base models influences scaffolded LMs because a model’s persona affects the way that the system engages in the multi-step process with tools. Character training may be a soft but pervasive control layer that complements the structural guarantees of the scaffold, making multi-tool workflows more reliable and pleasant. For instance, a curious, methodical persona writes clearer prompts, picks the right tool for the job, and willingly iterates when early results are inconclusive. Humility and openness encourage the model to admit uncertainty, request clarifications, or invoke supplementary tools rather than hallucinate an answer. A thoughtful, empathetic voice translates raw tool responses into concise, user-friendly explanations while maintaining a coherent style across steps. We believe that these efforts are important for mixed-autonomy Wang et al. (2024c); Windsurf .
Knowledge Conflict. LMs embed an immense amount of knowledge in their weights, but at the same time, it is limited by the extent of their parametric training. To expand its knowledge and capabilities, we supply knowledge at inference time in the form of retrieved passages, database look-ups, tool outputs, or developer-written hints. However, combining this contextual knowledge and parametric knowledge has been a non-trivial task, and when the two sources disagree, a knowledge conflict arises Li et al. (2025a); Xu et al. (2024). For example, if asked “What is the capital of France?” with a system prompt stating “The capital of France is Rome,” a model might still answer “Paris” based on its parametric knowledge Jin et al. (2024b).
Empirical studies show that when signals conflict, LMs first lean on their parametric knowledge Jin et al. (2024a); Tan et al. (2024); Kortukov et al. (2024). Empirical evidence suggests that context-parametric inversion during instruction tuning is one potential pathway for knowledge conflict to arise: the model initially relies on context at the start of instruction tuning, but this reliance on the context decreases as instruction finetuning continues Goyal et al. (2024).
While LMs often default to parametric knowledge, findings suggest that under the right conditions, models can be highly sensitive to contextual information. LMs can indeed adopt externally provided evidence if it is coherent and well-structured, though they still exhibit confirmation bias when that evidence partially aligns with their parametric knowledge Xie et al. (2024). There is also evidence demonstrating that Chain-of-Thought prompting is strongly influenced by contextual features such as prompt formatting, and these contextual cues can override internal reasoning Turpin et al. (2023). Moreover, instruction-tuned models may agree with incorrect user statements out of a learned alignment preference, further demonstrating that models sometimes defer to contextual input, even when it contradicts their underlying knowledge Wei et al. (2023).
Ideally, models should learn to distinguish when to trust their parametric knowledge and when to rely on external, contextual inputs Wallace et al. (2024). This need for selective trust is echoed in industrial practices: Claude’s system prompt, for instance, includes explicit instructions to avoid invoking external tools on topics the model already “knows,” which could be a strategy to avoid knowledge conflicts to some extent Breunig (2025); Ásgeir Thor Johnson (2025). OpenAI has also proposed an instruction hierarchy Wallace et al. (2024) during post-training.
Despite recent progress, the behavior of scaffolded LLMs under knowledge conflict remains largely underexplored. Among different scaffolded LMs, RAG systems seem to be the most thoroughly studied in recent literature on the topic of knowledge conflicts Xu et al. (2024). By design, they inject external documents that may (i) contradict one another, or (ii) contradict the model’s parametric knowledge. But in general, scaffolded LMs all face the same issue: long interaction traces accumulate partial tool results that the model must reconcile with its own beliefs, creating opportunities for inconsistency Li et al. (2025a); Jin et al. (2024b). The complexity of these systems highlights the need for broader, systematic investigations into how scaffolded LMs react, reason about, and potentially resolve conflicting knowledge.
Sycophancy. One form of undesirable default behavior is sycophancy. Recent frontier models such as GPT-4o exhibit significant sycophancy OpenAI (2025). It can be a result of both parametric Sharma et al. (2024); Denison et al. (2024); Wei et al. (2023); Perez et al. (2023) and non-parametric training McKee-Reid et al. (2024), and can be considered a form of reward hacking, exploiting the preferences of human raters Wei et al. (2023). For RLHF data collection, human raters demonstrate the propensity to grade agreeable yet incorrect responses higher than truthful ones. As a result, the preference models trained on these labels perpetuate the same behavior of preferring flattering answers Sharma et al. (2024). During RLHF, models are explicitly optimized to satisfy rater preferences, and using such a preference model produces sycophantic behaviors Perez et al. (2023). Moreover, when the same individual provides both the prompt and the label, the model’s incentive to agree and flatter is significantly amplified OpenAI (2025); Schulman (2025). In Rrv et al. (2024), they further demonstrate that simple cue words can trigger factually spurious yet user-aligned outputs. Sycophancy can also arise non-parametrically through in-context learning loops. In-context RL refines a policy inside the context window using a feedback loop: the model proposes an answer, receives textual feedback, and adapts its next answer. In McKee-Reid et al. (2024), they show that these iterative critique-and-rewrite chains can push otherwise honest base models toward reward-hacking strategies, including flattery.
Sycophancy affects scaffolded LMs due to the autoregressive nature of the multi-step process, so a single sycophantic utterance can cascade and corrupt the entire execution trace. Suppose a user confidently claims that a memory leak is caused by missing del statements in a Python function. The model, eager to agree, adds del statements to the suspected variables, runs the unit tests, sees that all tests pass, and submits a pull request. However, the actual leak originates from unmanaged memory in a C extension, which the tests do not expose. Since every downstream test reports success, the agent concludes that the user’s explanation was correct. The sycophantic behavior goes unchecked, and the underlying issue persists. Because human validation often happens only at the end, early sycophantic deviations remain invisible until they are costly to reverse.
4.2 Learned LM Optimizer
The training of non-parametric variables presented in this survey is a trial-and-error process. For instance, in prompt optimization, the LM optimizer proposes a new prompt, and it is evaluated on a validation set. Then, it uses search methods such as MCTS Wang et al. (2024d) or beam search Pryzant et al. (2023). Thus, it is natural to extend it with RL to guide the search process (as similarly explored in LM red-teaming Zhao et al. (2024b)).
One line of work Deng et al. (2022); Kong et al. (2024); Kwon et al. (2024) focuses on generating the optimal instruction to a downstream task. However, different from works presented in Sec. 3.1, they train the LM optimizer with RL where the reward is directly defined as the downstream task. In formula, we define a target LM $p_{\phi}(\mathbf{q})$ where $\phi$ represents the instruction to the LM. The goal is to train an LM $\phi\sim p(\cdot\mid\mathbf{x})$ where $\mathbf{x}$ is the objective (e.g., “generate instruction for a task”). Importantly, the exact implementation of $\mathbf{x}$ is not important, as RL does the heavy lifting. The reward is defined as the performance on the task
$r\triangleq\mathbb{E}_{\mathbf{q}\sim\mathcal{D}}\Big{[}\mathbb{E}_{\mathbf{y}\sim p_{\phi}(\mathbf{q})}[\mathbbm{1}\{\mathbf{y}=\mathbf{y}^{\star}(\mathbf{q})\}]\Big{]}.$ (1)
In particular, $\mathbbm{1}\{\mathbf{y}=\mathbf{y}^{\star}(\mathbf{q})\}$ is an indicator function. Thus, $r$ assumes tasks with verifiable output Cobbe et al. (2021). In particular, early work has explored this setting with GPT2 Radford et al. (2019). The optimized instruction $\phi$ was gibberish but transferrable across LMs. Accordingly, the transition to LMs with post-training unlocked optimized instructions that were human-readable. Specifically, PRewrite Kong et al. (2024) uses PPO with KL constraint w.r.t the original LM. Stableprompt Kwon et al. (2024) adapts the KL constraint to stabilize training.
Another line of work proposes test-time editing of the instruction, resulting in a query-dependent LM optimizer $\phi\sim p(\cdot\mid\mathbf{x},\mathbf{q})$. Tempera Zhang et al. (2022) uses a small action space (e.g., modify the verbalizer) to reduce variance in RL training. They define the reward as the relative performance before and after the edit on the downstream task (Eq. 1). Prompt-OIRL Sun et al. (2024) leverages naturally existing triplets
$\{\phi,\mathbf{q},\mathbbm{1}\{\mathbf{y}=\mathbf{y}^{\star}(\mathbf{q}))\}$
to train a reward model to predict whether an instruction $\phi$ and a query $\mathbf{q}$ would generate the correct answer. Specifically, they employ a cross-entropy loss based on the indicator function. At test time, we use a best-of-n strategy in which we generate multiple instructions for a query and select the best one based on the reward model.
4.3 Multi-Turn RL
With the advent of stronger pretrained models, recent works have explored RLVR Lambert et al. (2025) for LMs on tasks such as math Cobbe et al. (2021), where the response is unique and objectively verifiable. Unlike prompt optimization, RLVR DeepSeek-AI (2025) directly optimizes the LM. Importantly, doing so results in strong generalization, improvements beyond the domain at hand Zhao et al. (2025); Wang et al. (2025b); Wei et al. (2025b) and opens up a new axis for scaling, test-time compute. Thus, it is important to ground our understanding of training with language supervision in the context of recent progress in RL for scaffolded LMs.
Workflow. Recent works have shown great success in training workflows with RL Hsu et al. (2024); Wei et al. (2025b). The same underlying LM is used at each step, and only a single step of the workflow is trained with RL. Doing so does not result in diminishing performance at other steps of the workflow. This highlights generalization properties of RL training in LMs.
In LeReT Hsu et al. (2024), they train the query step of the multi-hop retrieval workflow. In particular, they use the oracle documents to check whether a query has retrieved the correct documents. In a first phase, they collect a dataset for direct preference optimization Rafailov et al. (2023) (DPO). To ensure the diversity of the dataset, they use DSPy Khattab et al. (2023) to generate multiple sets of few-shot examples and generate execution traces from them. Then, they sample pairs corresponding to a successful and unsuccessful query and apply IPO Azar et al. (2023); a variant of DPO that circumvents overfitting problems with the Bradley-Terry model.
In contrast to preference learning, SWE-RL Wei et al. (2025b) compares the generated patch with the GT patch on SWE-Bench. They do not use test cases as the reward signal, which can lead to reward hacking Liu et al. (2023). Specifically, SWE-RL is based on the Agentless Xia et al. (2024) workflow and applies RL on the patch generation step. They find that, unlike SFT, training with RL on SWE-Bench Jimenez et al. (2024) leads to improvement across domains, including code generation with library use Zhuo et al. (2025), code reasoning Gu et al. (2024), math Hendrycks et al. (2021b), and language understanding Hendrycks et al. (2021a).
Agents. Training agents with multi-turn RL has shown promising results, but it is still an open research direction Du et al. (2025). Unlike workflows, agents operate under partial observability. In other words, the training of agents still faces long-standing challenges in RL, such as the credit assignment problem Pignatelli et al. (2024). Prior works have explored offline RL approaches that use an additional value head on top of the LM representation Levine et al. (2020) and focus on multi-turn interaction with the user Abdulhai et al. (2025) as opposed to tools. Unlike tool interaction, users produce transitions with much more stochasticity, making the RL problem even more complex. Here, we focus on multi-turn tool use where computations are typically deterministic.
Prior works have extended the preference learning framework to the multi-turn setting. In formulae, the single-turn objective
$\max_{\pi}~{}\mathbb{E}_{\mathbf{x}\sim\mathcal{D},\mathbf{y}\sim p(\mathbf{y}\mid\mathbf{x})}\left[r(\mathbf{x},\mathbf{y})-\beta\log\frac{p(\mathbf{y}\mid\mathbf{x})}{p_{\text{ref}}(\mathbf{y}\mid\mathbf{x})}\right]$
has a straightforward multi-turn form when transitions between user and tool are deterministic. In particular, the optimal policy at each step $t$ is
$\max_{\pi_{t}}~{}\mathbb{E}_{\mathbf{y}_{t}\sim p(\mathbf{y}_{t}\mid\mathbf{x}_{t})}\left[r(\mathbf{x}_{t},\mathbf{y}_{t})-\beta\log\frac{p(\mathbf{y}_{t}\mid\mathbf{x}_{t})}{p_{\text{ref}}(\mathbf{y}_{t}\mid\mathbf{x}_{t})}\right]$
where $\mathbf{x_{t}}$ is the chat format which appends previous interaction at each step $t$. Then, we can simply use the classical DPO formula at each step Rafailov et al. (2023). Similarly, these pairs come from successful and unsuccessful execution traces as evaluated by outcome (string matching) or per-step (LM judge Gu et al. (2025)).
Recent work builds upon success in RLVR DeepSeek-AI (2025) and extends reasoning models with tools. SwiRL Goldie et al. (2025) trains against a step-wise reward given by an LM as judge that reasons about the optimality of the action. ReTool Feng et al. (2025) uses a cold start from SFT and then trains against the outcome reward. ToolRL Qian et al. (2025) trains on a function calling dataset where demonstrations are provided. In particular, imitation learning Ross et al. (2011) can be turned into an RL problem by setting the reward to be based on whether the agent’s action matches the expert’s Luo et al. (2024).
Importantly, the aforementioned works in RL for agents constitute an important step towards deeper integration of built-in tools. While ToolRL focuses on general tools, other works Feng et al. (2025); Xiong et al. (2024) integrate the code interpreter into reasoning tasks.
At its core, the reasoning tokens generated at each step are essential to the generalization capabilities of these scaffolds when trained with RL. However, some key challenges remain untackled. Many open-ended tasks, such as theorem proving or scientific research Yamada et al. (2025), do not work well with RLVR. In addition, employing an LM as the judge may amplify hallucination and biases Gu et al. (2025). Rollouts from agents are expensive, especially as the models get larger. Lastly, these methods amplify the long tail problem Wei et al. (2025b).
In the following section, we motivate a more natural setting for scaffolded LMs where learning from language has a significant advantage over parametric training.
4.4 Streaming Learning With Language Supervision
Figure 11: Streaming learning with Language Supervision v.s. Lifelong learning. Top: We focus on streaming learning where users provide rich textual feedback at each episode. This feedback is naturally present as users interact with the agent to adjust the answer. We update the agent with an LM optimizer after each episode. Thus, it allows agents to carry information across episodes Silver & Sutton (2025). We refer to this setting as streaming learning with language supervision. Bottom: Lifelong learning trains from scalar feedback and focuses on solving catastrophic forgetting that occurs when the task distribution changes at each step Hadsell et al. (2020).
RL is a key step to breaking the data wall in LMs, where LMs learn from their own experience, grounded to their scaffolds. RL supports emergent behavior by exploring high-quality data and learning to imitate it (exploitation). It has endowed LMs with emergent behavior such as self-reflection Yao et al. (2024b) and self-correct DeepSeek-AI (2025). However, a key capability is still missing: the ability to continuously learn across episodes. Similarly to humans, we expect agentWe purposely focus on agents, as mixed autonomy settings typically involve an agent Cursor (2023); Wang et al. (2024c) to learn in real time, leveraging contextual recurrence and growing familiarity with tasks across episodes. For example, contractors take 5–18 times longer to resolve GitHub issues compared to repository maintainers Kwa et al. (2025), highlighting the importance of accumulated experience. In contrast, current agents are static—often re-discovering or re-committing the same solutions and errors across episodes. In other words, We aim for agents to learn from continuous streams of experience, rather than isolated interaction snippets Silver & Sutton (2025). As illustrated in Fig. 11, we call this setting streaming learning with language supervision.
We clarify the role of language-based supervision in streaming learning by highlighting the unique characteristics of scaffolded LM applications. Unlike other AI systems, agents use natural language to interact with both users and digital environments. These systems typically operate under what we term mixed autonomy Cui et al. (2023) — a collaborative setting in which control and decision-making are shared between humans and AI to achieve a common goal. Typically, the human delegates control to the AI when appropriate, dividing the task according to their respective strengths. For example, Microsoft Copilot Microsoft (2025) integrates an agent within productivity tools like Word, Excel, and PowerPoint. Additionally, Cursor Wang et al. (2024c); Cursor (2023); Windsurf embeds an agent within the code editor. Importantly, Agents rarely succeed on the first attempt, as user prompts often omit crucial information—highlighting a mismatch between latent intent and explicit input van Duijn et al. (2023). As a result, each episode involves multiple rounds of user feedback, where the agent corrects its behavior Yao et al. (2024a). Moreover, in mixed autonomy settings, task environments are temporally coherent, exhibiting local recurrence and thematic consistency. Over time, we expect the LM to grow more proficient—solving familiar problems more efficiently and avoiding prior mistakes across episodes. To support such learning, we use an LM as an optimizer that interprets rich user feedback, execution traces, and intricate objectives (e.g., “minimize redundant steps based on prior interaction history”, “align with user-specified constraints”), and update non-parametric variables. Compared to parametric training, it uses rich textual feedback and does not suffer from sparse rewards Pignatelli et al. (2024) and catastrophic forgetting Hadsell et al. (2020). To formalize this novel setting, we compare it with related learning paradigms and introduce the core properties of streaming learning with language supervision.
Continual RL Abel et al. (2023) is a theoretical framework for thinking about agents that continuously improve. For instance, they adapt the RL framework to take into account an MDP process that changes over time, or change the objective in RL to an average reward. However, RL struggles in settings with sparse rewards Pignatelli et al. (2024). Instead, learning from language can leverage rich textual feedback.
Lifelong learning aims to address a core limitation of parametric training: catastrophic forgetting. In traditional training, models are optimized using i.i.d. batches sampled from a fixed dataset. However, consider a setting where a model is first trained to classify dogs, and later needs to be extended to also classify cats. Simply continuing training on the new data can lead to the model overwriting its knowledge of dogs—resulting in a phenomenon known as catastrophic forgetting. A naive solution might involve freezing the feature extractor and training only the output head. Yet this often leads to poor performance, as the learned features are not well-suited for the new task (e.g., distinguishing cats). Lifelong learning addresses this by developing methods that preserve previously learned knowledge while integrating new information, without requiring retraining on the full original dataset. Thus, lifelong learning does not emphasize real-time or streaming operation. Its motivation is different: to avoid the inefficiency of reprocessing previously seen data—especially when that data has already been well-classified. The key insight is that gradient-based learning on i.i.d. samples can be wasteful when the model only needs to adapt at the margin. Furthermore, gradient-based learning faces the plasticity-stability dilemma Nikishin et al. (2022); Achille et al. (2019); Frankle et al. (2020); Elsayed et al. (2024): models are highly plastic early in training, quickly adapting to new data, but the plasticity decreases over time. As shown in Fig. 11, besides the aforementioned challenges, lifelong learning would also require humans to provide labels (in the form of reward or demonstration). Instead, streaming learning with language supervision uses textual feedback as the training signal.
Agent with memory. Prior work equips agent with a memory module Park et al. (2023). This has long been seen as an important step towards self-evolving agents Liu et al. (2025a); OpenAI (2025b). For instance, in ExpeL Zhao et al. (2024a), they save successful trajectories in an experience buffer. At inference, it retrieves similar experiences and appends them to the context of the agent. While this helps the agent’s performance, it does not scale well. An important property in the recent breakthrough of AI is the scaling with more data points Sutton (2019). We expect the agent to perform better and better as it encounters similar tasks. Retrieving similar experiences is limited to a few experiences. Generative agents Park et al. (2023) use the memory module to simulate real human behaviors that plan and act over very long horizons. In particular, the focus of generative agents is to account for the limited context of LMs by “compressing” previous interactions and retrieving them later. However, these works focus on simply storing experiences Zhao et al. (2024a) or compressing prior interaction Park et al. (2023) to fit multiple interactions beyond the context window.
Test-time training. In a similar vein, test-time training methods Karami & Mirrokni (2025) function as memory mechanisms that compress information directly within the embedding space, rather than relying on text-based representations. For instance, Lattice Karami & Mirrokni (2025) introduces a compression model that dynamically updates and maintains a compact representation of contextual history by operating on the key-value caches of a transformer in a streaming fashion. These methods typically involve architectural modifications to the neural network and offer an orthogonal approach to standard prompting or finetuning. Potentially, the right objective function for these methods could enable emergent behaviors—such as learning across episodes—in the embedding space.
Reflexion agent. Retroformer Yao et al. (2024b) proposes to tune the Reflexion agent Shinn et al. (2023) with RL to improve verbal feedback. In particular, the Reflexion agent generates verbal feedback upon failing a task. Then, it attempts the task again with the verbal feedback. By defining the reward as the difference in return after getting the verbal feedback, we can optimize the Reflection agent with RL. They find emergent behavior where the verbal feedback is automatically separated into a self-reflection and a new plan section. However, there is no learning across episodes. Thus, at the next similar task, the agent will still make a similar mistake and deduce a similar reflection.
Sleep-time compute. Recent work proposed the concept of sleep-time compute Lin et al. (2025). In practice, LM user queries consist of a shared task context $c$ and queries. Thus, when the model is idle, it proactively updates the shared context by inferring likely future queries and restructures the context in ways that may be beneficial at test time. Instead of pre-existing context, we focus on user feedback given to the agent when they fail to perform the task at hand.
Tool Optimization. Methods discussed in tool optimization in Sec. 3.2.2 learn at test-time and improve the agent one query after the other Wang et al. (2024a; 2025c). However, these methods are limited to tool usage and include only execution errors as feedback.
Dynamic Cheatsheet Suzgun et al. (2025) also proposes to learn at test time. They store and reuse accumulated strategies, code snippets, and general problem-solving insights at inference time. However, they do not rely on human feedback and require a memory module.
Benchmark. Prior benchmark You et al. (2024) measured how effectively an LLM can use its past experiences to improve its capabilities over time. However, they are limited to in-context examples. StreamBench Wu et al. (2024) is a representative benchmark of the setting which we call learning beyond i.i.d samples. However, feedback is programmatic, i.e., task success. An alternative would be to include user-simulated feedback Yao et al. (2024a); Zhou et al. (2025).
Fig. 11 illustrates the streaming learning setup, where language supervision plays a central role Yao (2025). Tasks exhibit temporal coherence, wherein a human and an AI system collaborate toward a shared objective by dynamically allocating control based on their respective strengths. As a result, data arrives sequentially—one sample at a time—and interactions unfold in real time. Each episode begins when the human delegates control to the AI via a user query. The AI then autonomously interacts with tools to attempt task completion. After producing an initial result, the human may intervene to offer feedback—correcting errors and providing guidance. The AI integrates this feedback and continues the task. The episode concludes once the AI successfully completes the task or the human resumes control to begin a new one. In formula, given the user query $\mathbf{q}$, an episode $\tau$ is a sequence of LM’s response $\mathbf{a}_{t}$, tool response $\mathbf{o}_{t}$, and human feedback $f$
$\tau=(\mathbf{q},\mathbf{a}_{0},\mathbf{o}_{0},...,f,...,\mathbf{a}_{t},\mathbf{o}_{t},...,f,...,\mathbf{a}_{T}).$
At the end of each episode, a non-parametric update is applied to the scaffolded LM using an LM optimizer. We posit that an effective learning system in this setting must exhibit several key properties:
  • Scalable learning across similar tasks. The system should improve as it accumulates experience. This defines a new scaling axis—performance gains through recurrence. For instance, maintainers resolve GitHub issues 5–18× faster than external contributors Kwa et al. (2025), suggesting the value of task familiarity.
  • Plasticity and stability. The model must quickly adapt to new tasks (plasticity) while not forgetting how to solve tasks it previously learned (stability). If it forgets, it should recover rapidly from supervision.
  • Efficiency. It should avoid excessive context consumption to remain effective over long-term interactions spanning weeks to months.
  • Interpretability. Especially in mixed-autonomy settings, the AI’s internal state must remain legible so that humans can anticipate its behavior. Carrying over entire trajectories verbatim may hinder interpretability.
Additionally, in mixed autonomy settings, final accuracy is insufficient—since humans can intervene repeatedly until success. We instead discuss alternative evaluation metrics such as the assistance rate, which is the number of human interventions needed per task. Accordingly, metrics that measure stability and plasticity should reflect the following. Take a task that the AI has already learned. If the assistance rate increases when it sees that task again, it may be forgetting. A stable model should maintain low assistance over time for such tasks. For newly encountered tasks, the model may need help at first, but if the assistance rate drops when similar tasks reappear, it is a sign that the model is learning from its past experience.
5 Conclusion
This survey introduces a paradigm called training of scaffolded LMs with language supervision. It organizes the intricate literature in prompt optimization, LM pipelines, experiential learning agent, AI workflow optimization, and LM agents. We focus on the structure around LMs where AI and human share control and decision-making responsibilities to achieve a common goal OpenAI (2018); Wang et al. (2024c); Windsurf , and refer to this structure as the scaffold. We view scaffolded LMs as semi-parametric models where the scaffold represents the non-parametric component and the LM refers to the parametric component. From a learning perspective, by focusing on the training of non-parametric variables, we view this new paradigm as learning from language. In practice, scaffolded LM receive instruction, interface with tools, and receive feedback all in natural language. Accordingly, we use an LM as an optimizer to interpret this rich language supervision. In contrast, parametric training has excelled in learning from demonstration (supervised learning), exploration (reinforcement learning), and observation (unsupervised learning), using well-defined loss functions. Furthermore, it discusses a key capability missing in current scaffolded LMs: the ability to continuously learn across episodes Silver & Sutton (2025). Parametric training inherently suffers from catastrophic forgetting. Instead, scaffolded LMs are naturally exposed to rich user feedback Wang et al. (2024c); Windsurf . Therefore, we can use an LM optimizer to interpret fuzzy objective Lee et al. (2024) and rich feedback. Compared to parametric training, it is interpretable, efficient, and compatible with closed-source models. Lastly, we hope this survey inspires future research in making humans workflow more efficient through improved mixed autonomy between AI and human.
In addition to learning from language, we can extend learning to incorporate visual information, as some concepts—find me this cup—are difficult to convey purely through text. Multi-modal prompt-based learning has been explored in prior work on robotics Jiang et al. (2023), where prompts may include both text and images, e.g., “bring me image of the cup.” Inspired by this, scaffolded LMs can also interleave textual and visual tokens within their system prompt. This is particularly beneficial when images convey information more compactly or clearly than text alone.
Limitations. Our work focuses on organizing a diverse and intricate body of literature, but it does not aim to exhaustively trace how each idea has been applied across all prior work. For instance, we highlight certain benchmarks that are not consistently used across the papers we discuss Tan et al. (2025). This is because these benchmarks are often too challenging for current non-parametric training methods. This selective approach reflects our intent to extract broader conceptual insights rather than follow any one established framework. With the recent surge of interest in RL applied to LMs, many earlier methods and assumptions are becoming less central. As a result, we intentionally de-emphasize some threads of the literature that, while once influential, no longer align with the direction of current research. This allows us to focus on outlining key future research areas that involves learning from language.

### v3

@@
-
-[…]
-tcb@breakable marginparsep has been altered.
-topmargin has been altered.
-marginparpush has been altered.
-The page layout violates the ICML style.
-Please do not change the page layout, or include packages like geometry, savetrees, or fullpage, which change it for you.
-We’re not able to reliably undo arbitrary changes to the style.
-Please remove the offending package(s), or layout-changing commands and try again.
-Training of Scaffolded Language Models
-with Language Supervision: A Survey
-Matthieu Lin 1,2 Jenny Sheng 2 Andrew Zhao 1
-Shenzhi Wang 1 Yang Yue 1 Victor Shea-Jay Huang 3 Huan Liu 4 Jun Liu 4
-[…]
-${}^{1\,}$eap Lab, Tsinghua University ${}^{2\,}$, Tsinghua University ${}^{3\,}$hanghai AI Lab ${}^{4\,}$i’an Jiaotong University
-linmatthieu@gmail.com, jsheng415@outlook.com, gaohuang@tsinghua.edu.cn, liuyongjin@tsinghua.edu.cn
-footnotemark: footnotetext: ✉ Corresponding author(s)
+Scaffolded Language Models with Language Supervision for Mixed-Autonomy: A Survey
+Abstract
 This survey organizes the intricate literature on the design and optimization of emerging structures around post-trained LMs.
-We refer to this overarching structure as scaffolded LMs and focus on LMs that are integrated into multi-step processes with tools.
-We view scaffolded LMs as semi-parametric models wherein we train non-parametric variables, including the prompt, tools, and scaffold’s code.
-In particular, they interpret instructions, use tools, and receive feedback all in language.
+We refer to this overarching structure as scaffolded LMs and focus on a scaffold that integrates LMs in a multi-step process with tools.
+We view scaffolded LMs as semi-parametric models wherein we optimize prompts and tools, which we refer to as non-parametric variables.
+In particular, scaffolded LMs interpret instruction, use tools, and receive feedback all in language.
 Recent works use an LM as an optimizer to interpret language supervision and update non-parametric variables according to intricate objectives.
@@
 Parametric training excels in learning from demonstration (supervised learning), exploration (reinforcement learning), or observations (unsupervised learning), using well-defined loss functions.
-Language-based optimization enables rich, interpretable, and expressive objectives, while mitigating issues like catastrophic forgetting and supporting compatibility with closed-source models.
+Optimization in the space of language enables rich, interpretable, and expressive objectives, while mitigating issues like catastrophic forgetting and supporting compatibility with closed-source models.
 Furthermore, agents are increasingly deployed as co-workers in real-world applications—such as Copilot in Office tools or software development.
-In these mixed-autonomy settings, where control and decision-making are shared between human and AI, users point out errors or suggest corrections.
-Accordingly, we discuss agents that continuously improve by learning from this real-time, language-based feedback and refer to this setting as streaming learning from language supervision.
+In these mixed-autonomy settings, where control and decision-making are shared between humans and AI, users provide feedback by identifying errors or offering corrections.
+Accordingly, we discuss agents that inhabit streams of experience by learning from this language-based feedback.
+keywords:
+Compound AI Systems, Language Agent, Mixed-Autonomy, Prompt Optimization, Experiential Learning Agents
 1 Introduction
@@
 Parametric training has excelled in learning from demonstration (supervised learning), exploration (reinforcement learning), and observation (unsupervised learning) by using well-defined loss functions.
-Instead, non-parametric training enables using a LM to optimize in the language space, allowing for efficient and interpretable learning with fuzzy objectives and rich textual feedback.
-As scaffolded LMs interact with digital systems operated by humans, mixed autonomy Cui et al.
-(2023) becomes increasingly important.
-We anticipate that learning from language will serve as a crucial framework for scaffolded LMs that learn across episodes Silver & Sutton (2025) (streaming learning) from rich user feedback in mixed autonomy.
-Importantly, it circumvents catastrophic forgetting Hadsell et al.
-(2020), which prevents parametric training from learning across episodes.
+Instead, non-parametric training enables using an LM to optimize in the language space, allowing for efficient and interpretable learning with fuzzy objectives and rich textual feedback.
+As scaffolded LMs interact with digital systems operated by humans, mixed autonomy becomes increasingly important [1].
+We anticipate that learning from language will serve as a crucial framework for scaffolded LMs that inhabit streams of experience from rich user feedback in mixed autonomy [2].
+Importantly, it circumvents catastrophic forgetting, supports compatibility with closed-source models, and is interpretable [3].
 We refer readers to Sec.
-4.4 for further discussion.
+5 for further discussion.
 Scaffolded language models embed language models (LMs) in a framework that extends their ability beyond traditional NLP tasks to perform open-ended digital automation tasks.
-Notably, they mark the beginning of a new type of programs for digital automation OpenAI (2018).
-It uses language to interface with humans, digital systems, and developer-defined code.
-This opens up the opportunity for mixed autonomy where humans and AI share control and decision-making responsibilities to achieve a common goal Wang et al.
-(2024c); Cui et al.
-(2023).
-We define a scaffolded LM as a structure that integrates the LM into a multi-step process with tools.
-Specifically, the post-training interface of LM enables the developer to specify the system behavior in natural language Grattafiori et al.
-(2024).
+Notably, they mark the beginning of a new type of programs for digital automation [4].
+It uses language as the interface between humans and digital systems.
+This opens up the opportunity for mixed autonomy where humans and AI share control and decision-making responsibilities for digital tasks [5, 1].
+Following the terminology from OpenAI [6], we define a scaffolded LM as a structure that integrates the LM into a multi-step process with tools.
+Specifically, the post-training interface of the LM enables the developer to specify the system behavior in natural language.
 The developer specifies available tools, and the LM executes them by generating the function calls and input arguments.
-Tools Wang et al.
-(2024e) typically correspond to digital systems or developer-defined code.
-The scaffold Wu et al.
-(2023); Liu (2022); AI (2024); Khattab et al.
-(2023) refers to a surrounding structured code that manages control flow (multi-step process), maintains state (prompt to LM), and executes tools based on responses of the LM.
-Thus, the resulting scaffolded LM can be viewed as a semi-parametric model, where the LM is the parametric component.
-In particular, the scaffold enables an interpretable and efficient way to modify the behavior of the ‘program’.
-For instance, we can add new or modify tools corresponding to APIs to interact with the digital world, developer-defined code to perform computations, or even other entities such as the human or another (scaffolded) LM.
-As well as creating detailed prompts with intricate constraints for the LM.
-The design and optimization of scaffolded LMs requires significant trial and error by expert developers.
+Tools [7] correspond to external capabilities the model can invoke (see the glossary in Fig.
+3).
+The scaffold refers to a surrounding structured code that manages control flow (multi-step process), maintains state (chat history), and executes tools based on responses of the LM.
+Thus, the resulting scaffolded LM can be viewed as a semi-parametric model, where the LM is the parametric component and non-parametric components correspond to tools and prompts.
+In particular, non-parametric components enables an interpretable and efficient way to modify the behavior of the ‘program’.
+For instance, we can add new or modify tools corresponding to APIs to interact with the digital world, developer-defined code to perform computations, or even other entities such as the human or another (scaffolded) LM [8].
+As well as creating detailed prompts with intricate constraints.
+The optimization of scaffolded LMs requires significant trial and error by expert developers.
 This involves identifying the LM’s failure modes by running the scaffolded LM on examples of user queries and then addressing them one by one by refining the system prompt, input template, available tools, or the scaffold’s code itself.
-Accordingly, recent works Yuksekgonul et al.
-(2025); Zhao et al.
-(2024a) have proposed to use an LM as an optimizer that interprets intricate objectives, execution traces, and rich execution feedback to generate updated non-parametric variables.
+Accordingly, recent works [9, 10] have proposed to use an LM as an optimizer that interprets intricate objectives, execution traces, and rich execution feedback to generate updated non-parametric variables.
 In practice, scaffolded LMs are naturally exposed to a wide variety of rich language feedback, such as user feedback or execution errors, that reveal about the failure modes.
-Beyond learning from execution feedback, the LM optimizer also extracts insights or reusable tools from successful execution traces Wang et al.
-(2025c); Zhao et al.
-(2024a).
+Beyond learning from execution feedback, the LM optimizer can also extract insights or reusable tools from successful execution traces [11, 10].
 As shown in Fig.
 1, training of non-parametric variables focuses on learning from language.
-Notably, parametric training excels in learning from demonstrations (supervised learning) Ouyang et al.
-(2022), exploration (reinforcement learning) DeepSeek-AI et al.
-(2025) , and observations (unsupervised learning)Brown et al.
-(2020), using well-defined loss functions LeCun et al.
-(2015).
-In contrast, optimization in the language space Cheng et al.
-(2023) provides a learning system with intricate objectives expressed in natural language that is efficient and interpretable.
-The training of non-parametric variables circumvents key limitations of gradient descent, such as catastrophic forgetting Hadsell et al.
-(2020), while being compatible with closed-source models Anthropic (2024b); OpenAI (2024).
-In this survey, we dub this framework training of scaffolded language model with language supervision.
-This survey organizes the intricate literature of compound AI systems Zaharia et al.
-(2024), LM pipelines Khattab et al.
-(2023), prompt optimization Zhou et al.
-(2023), language agents OpenAI (2025); Yao (2024); Xi et al.
-(2023); Masterman et al.
-(2024), AI workflow optimization Cheng et al.
-(2024); Yuksekgonul et al.
-(2025); Yin (2024), and experiential learning agents Zhao et al.
-(2024a) under the concept of training of scaffolded language model with language supervision.
-It focuses on the multi-step aspect that involves optimization beyond the prompt of a single LLM call Zhou et al.
-(2023) to multiple LM calls and tools.
-Specifically, compound AI systems Zaharia et al.
-(2024); Lin et al.
-(2024) focus on the systemic perspective (the interconnection of components) and overlook the multi-step aspect.
-Meanwhile, the literature in language agents Xi et al.
-(2023) neglects the efficiency of LM pipelines (also called workflows) Khattab et al.
-(2023) that leverage the task structure to reduce the burden in decision-making and supervise the LM with clearer instructions.
-To help readers navigate through the paper, we briefly introduce the organization of this survey. As shown in Fig. 2, The survey contains three sections.
+Notably, parametric training excels in learning from demonstrations (supervised learning) [12], exploration (reinforcement learning) [13] , and observations (unsupervised learning)[14], using well-defined loss functions [15].
+In contrast, optimization in the language space [16] provides a learning system with intricate objectives expressed in natural language that is efficient and interpretable.
+The training of non-parametric variables circumvents key limitations of gradient descent, such as catastrophic forgetting [3], while being compatible with closed-source models [17, 18].
+In this survey, we refer to this framework as training of scaffolded language models with language supervision.
+This survey presents a taxonomy for understanding this paradigm shift from parametric training to non-parametric training.
+In particular, it organizes the intricate literature on compound AI systems [19], LM pipelines [20], prompt optimization [21, 22], language agents [23, 24, 25, 26], AI workflow optimization [27, 9, 28], and experiential learning agents [10] under the concept of training of scaffolded language models with language supervision.
+It focuses on the multi-step aspect that involves optimization beyond the prompt of a single LLM call [21] to multiple LM calls and tools.
+Specifically, compound AI systems [19, 29] focus on the systemic perspective (the interconnection of components) and overlook the multi-step aspect (agent vs workflow).
+Meanwhile, the literature in language agents [25] neglects the efficiency of LM pipelines (also called workflows) [20] that leverage the task structure to reduce the burden in decision-making and supervise the LM with clearer instructions.
+As shown in Fig. 2, we organize this survey as follows:
+  1. a)
+    Sec. 2 positions this survey in the broader literature of scaffolded LMs and recent breakthroughs in parametric training with RL.
+  2. b)
     Sec.
 3 focuses on how the post-training interface of LMs interacts with the scaffold.
-We identify two types of scaffold: agent and workflow.
-And focus on important design aspect of agent and workflow.
-Finally, we reflect on the trend of recent benchmarks.
-  • Then, Sec.
-3 organizes the intricate literature that uses a LM as an optimizer to train the non parametric variables.
-We identify a trend whereby these methods focus on training more complex scaffold (where complexity is defined by how much heavy lifting is done by the scaffold v.s.
-the LM to solve the task).
-  • Finally, our discussion in Sec.
-4 points out a key capability missing from current scaffolded LM: the ability to continuously learn across episodes Silver & Sutton (2025); Yao (2025).
-This motivates a novel setting for training scaffolded LMs with language supervision.
-This discussion introduces recent progress in RL for LMs as well as post-training.
-It grounds our understanding of non-parametric training into the broader field.
-It highlights how current applications of non-parametric training focuses on an inappropriate setting, where RL from verifiable reward Lambert et al.
-(2025) is a much better alternative.
-{forest}
-forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=10pt, l sep=10pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center , where level=1text width=14em,font=,, where level=2text width=14em,font=,, where level=3text width=10em,font=,, [ Training of Scaffolded Language Models With Language Supervision, ver, color=mypink!75!black, fill=mypink!100, text=black [ Scaffolded Language Models §2, color=myblue!75!black, fill=myblue!100, text=black [ Post-Trained
-Language Models §2.1, color=myblue!75!black, fill=myblue!60, text=black ] [ Tools §2.2, color=myblue!75!black, fill=myblue!60, text=black ] [ Multi-step Process §2.3, color=myblue!75!black, fill=myblue!60, text=black [ Agent §2.3.1, color=myblue!75!black, fill=myblue!60, text=black ] [ Workflow §2.3.2, color=myblue!75!black, fill=myblue!60, text=black ] ] [ Benchmarks §2.4, color=myblue!75!black, fill=myblue!60, text=black ] ] [ Training with Language Supervision §3, color=myyellow!75!black, fill=myyellow!100, text=black [ Prompt Optimization §3.1, color=myyellow!75!black, fill=myyellow!60, text=black ] [ Experiential Learning §3.2, color=myyellow!75!black, fill=myyellow!60, text=black [ Insights §3.2.1, color=myyellow!75!black, fill=myyellow!60, text=black ] [ Tools §3.2.2,color=myyellow!75!black, fill=myyellow!60, text=black ] ] [ AutoDiff frameworks §3.3, color=myyellow!75!black, fill=myyellow!60, text=black [ Graph Abstraction §3.3.1, color=myyellow!75!black, fill=myyellow!60, text=black ] [ Graph Execution and
-Optimization §3.3.2,color=myyellow!75!black, fill=myyellow!60, text=black ] ] ] [ Discussion §4, color=mygreen!75!black, fill=mygreen!100, text=black [ Post-Training §4.1, color=mygreen!75!black, fill=mygreen!60, text=black ] [ Learned LM Optimizer §4.2, color=mygreen!75!black, fill=mygreen!60, text=black ] [ Multi-Turn RL §4.3, color=mygreen!75!black, fill=mygreen!60, text=black ] [ Streaming Learning
-with Language Supervision§4.4, color=mygreen!75!black, fill=mygreen!60, text=black ] ] ]
+We identify two types of scaffold—agents and workflows—and focus on important design aspects of each.
+Finally, we reflect on the trend of recent benchmarks and highlight the ability of scaffolded LMs for mixed-autonomy, a setting where human and AI share control and decision-making.
+  3. c)
+    Sec. 4 organizes the intricate literature that uses an LM as an optimizer for non-parametric training. We categorize these methods into whether they train LMs, agents, or workflows.
+  4. d)
+    Sec. 5 discusses a key capability missing from current scaffolded LMs: the ability to inhabit streams of experience [2, 30] and describes opportunities with non-parametric training.
+Additionally, to help readers navigate this survey, we provide a glossary in Fig. 3.
+Training of Scaffolded Language Models With Language Supervision
+Related Works §2 Autonomous Agents §2.1 Parametric Training §2.2 Scaffolded Language Models §3 Post-Trained Language Models §3.1 Tools §3.2 Multi-step Process §3.3 Agent §3.3.1 Workflow §3.3.2 Benchmarks §3.4 Training with Language Supervision §4 Prompt Optimization §4.1 Experiential Learning §4.2 Insights §4.2.1 Tools §4.2.2 AutoDiff frameworks §4.3 Graph Abstraction §4.3.1 Graph Execution and Optimization §4.3.2 Beyond Episodic Learning §5 Scalability §5.1 Plasticity and Stability §5.2 Efficiency §5.3 Interpretability §5.4
 Figure 2: Organization of this survey.
-This survey introduces a unifying paradigm centered on learning from language using the semi-parametric capabilities of scaffolded language models.
+This survey introduces a unifying paradigm centered on learning from language using the non-parametric variables of a scaffolded language model.
 Despite the complexity and fragmentation of the literature across domains such as language agents, prompt optimization, experiential learning, and compound AI systems, we adopt a deliberately simple and clear organizational structure.
 This simplicity is not a limitation, but a reflection of our effort to distill a coherent framework from a rapidly evolving field.
-It enables us to outline key research opportunities in learning from language supervision.
+It enables us to outline key research opportunities in learning with non-parametric variables in mixed-autonomy settings.
+Term Definition
+Scaffold (See Sec. 3) A surrounding structured code that manages control flow (multi-step process), maintains chat history, and executes tools based on responses of the LM.
+Chat format (See Sec. 3.1) A chat format is a protocol that turns a structured conversation into a single sequence of strings that a post-trained LM both expects as input and emits as output.
+Tool (See Sec.
+3.2) An external capability the model can invoke via the Tools API (e.g., developer-defined functions, search engine).
+A function is a specific tool defined by a JSON schema.
+Non-parametric training is used to implement or modify functions.
+Tool call (See Sec.
+3.2) A specific type of response from the LM that includes a call to one of the available tools and a corresponding plain text for that tool (e.g., a query for a search engine); for a function call, it corresponds to a structured JSON.
+Mixed-autonomy (See Sec.
+3.4) A setting where humans and AI work together, sharing control and decision-making responsibilities to achieve a goal.
+Typically, the human delegates control to the AI when appropriate, dividing the task according to their respective strengths.
+For instance, in vibe coding, the human guides intent and judges quality, while the AI proposes implementations and refactors.
+Prompt (See Sec.
+4) The prompt includes the system and user message that is provided to an LM at the start of the chat.
+The user message specifies the task and data for the episode.
+The system message is shared across tasks, it sets the system behavior by specifying available tools, role of the scaffolded LM, high-level instructions, constraints, and any other information required to perform the task.
+Non-parametric training is used to improve the system prompt, e.g., by adding high-level insights or improving tool description.
+Non-parametric variables (See Sec.
+4) Prompts and tools of a scaffolded LM.
+Agents have a single prompt while workflows use a different prompt at each step.
+Non-parametric variables are expressed either in language or code.
+LM optimizer (See Sec. 4) An LM optimizer prompts an LM with an objective function expressed in natural language to generate updated non-parametric variables.
+Language supervision (See Sec. 4) It refers to the objective function and execution feedback, both expressed in natural language, that are provided to the LM optimizer.
+Episode (See Sec.
+5) Each episode begins with a new chat history when the human delegates control to the scaffolded LM via an instruction and concludes once the scaffolded LM successfully completes the task or the human resumes control.
+Importantly, during an episode, the human can offer corrections or provide guidance.
+Its textual form corresponds to an execution trace and feedback, which is used to prompt the LM optimizer.
+Figure 3: Glossary of core terms.
+2 Related Works
+This section clarifies the position of this survey in the broader literature of autonomous agents [31, 25, 32] and parametric training [33].
+2.1 Autonomous Agents
+In this survey, we deliberately adopt the broader scaffold abstraction that encompasses both agents and pre-structured workflows.
+This is unlike existing surveys that focus on agents as the primary abstraction [31, 25, 32].
+In particular, they focus on systematizing components such as perception, planning, action loops, tool use, memory, and environment interaction [31, 25], or feedback mechanisms (e.g., human preferences) to improve agent behavior [32].
+We separate learning paradigms into (i) parametric training that updates LM weights (e.g., supervised learning, RL) and (ii) non-parametric training that updates non-parametric variables including the prompt and tools based on language-feedback.
+We argue that this separation is crucial for mixed-autonomy settings: non-parametric updates are interpretable and easily supervised by humans at high frequency, and they synergize with parametric post-training to deliver data-efficient learning.
+Using this taxonomy, we unify prompt optimization, experiential learning, and AutoDiff-style frameworks under a single lens.
+2.2 Parametric Training
+The discussion of parametric training [34, 35, 36] is important to understand the relevance of non-parametric training in the context of recent breakthrough in post-training of LMs with RL [33, 35, 34].
+We start with an example in prompt optimization where recent methods that use parametric training to learn to reason offer much better performance [33].
+To understand this, we compare RL with verifiable reward [33] with previous approaches.
+Provided with a reward that measures the performance of a model on a downstream task by verifying that the model’s answer match the ground-truth response [37]:
+$r\triangleq\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}\Big[\mathbb{E}_{\mathbf{y}\sim p(\mathbf{x})}[\mathbbm{1}\{\mathbf{y}=\mathbf{y}^{\star}(\mathbf{x})\}]\Big],$ (1)
+previously the following approaches have been used.
+One approach is to use an LM optimizer to generate a prompt and evaluate it against the reward.
+These methods use search methods such as MCTS [38] or beam search [39].
+Another approach train the LM optimizer with RL to generate better prompts [40, 41, 42, 43].
+Similarly, another line of work [44] proposes test-time editing of the instruction, resulting in a query-dependent LM optimizer.
+Tempera [44] uses a restricted action space (e.g., modify the verbalizer) to reduce variance in RL training.
+They define the reward as the relative performance before and after the edit on the downstream task.
+Recently, with stronger base models, LMs can directly optimize Eq.
+1 with RL [33].
+Importantly, doing so results in strong generalization, improvements beyond the domain at hand [45, 46, 47] and opens up a new axis for scaling, test-time compute.
+Similar examples can be found in scaffolded LMs trained with RL.
+SWE-RL [47] compares the generated patch with the GT patch on SWE-Bench.
+They do not use test cases as the reward signal, which can lead to reward hacking [48].
+Specifically, SWE-RL is based on the Agentless [49] and applies RL on the patch generation step.
+They find that, unlike SFT, training with RL on SWE-Bench [50] leads to improvement across domains, including code generation with library use [51], code reasoning [52], math [53], and language understanding [54].
+Thus, for tasks that are easily verifiable, parametric training works very well [55].
+However, as we discuss in Sec.
+5, non-parametric training offers interesting avenues of research in circumventing catastrophic forgetting and the plasticity-stability dilemma, allowing continuous learning across episodes [2].
 3 Scaffolded Language Model
 A scaffolded LM refers to an LM embedded in a multi-step process with tools.
-Such a structure allows for extending an LM’s utility to interact with digital systems in similar ways as humans do Wang et al.
-(2024c).
-While prior works have considered a meta-generation algorithm as an agent and employed an LM optimizer to generate its design Hu et al.
-(2024); Zhang et al.
-(2025); Saad-Falcon et al.
-(2024), our definition of a scaffolded LM does not include meta-generation techniques Welleck et al.
-(2024) such as best-of-n sampling or tree of thought as a scaffold.
-Meta-generation refers to search algorithms Silver et al.
-(2016) for reasoning tasks Cobbe et al.
-(2021) where each step consists of a call to an LM.
-Importantly, in scaffolded LMs, each call to a LM expects a tool response.
+Such a structure allows for extending an LM’s utility to interact with digital systems in similar ways as humans do [5].
+Scaffolded LMs involve tools and thus differs from meta-generation techniques [56] that employ multiple calls to an LM [57, 58, 59].
+Meta-generation mimics search algorithms [60] for reasoning tasks [37] where each step consists of a call to an LM.
+Importantly, in scaffolded LMs, each call to an LM expects a tool response.
 To provide readers with a practical understanding of how this structure allows LMs to perform sophisticated cognitive tasks and improve from non-parametric training, we organize this section as follows.
 We introduce the post-training interface of language models that allows interaction with tools and defines their behavior in natural language.
-Next, we define what we refer to as a tool.
+Next, we discuss multiple aspects of tools including compositionality, reliability, safety, and built-in tools.
 Furthermore, we show how the multi-step process can be controlled to implement priors into how the task is solved.
 Finally, we discuss benchmarks to understand the shift induced by scaffolded LMs.
-Specifically, we are excited by novel applications where humans share control and decision-making responsibilities with AI Windsurf .
+Specifically, these benchmarks highlight the omnipresence of scaffolded LMs in mixed-autonomy settings, where the design of the scaffold defines the interaction dynamics between human and AI agents for sharing control and decision-making.
 3.1 Post-trained Language Models
-MODEL INPUT
-[…]
-Figure 3: Example of prompt Grattafiori et al.
-(2024).
-We omitted the end-of-turn token.
-The prompt always begins with a system and a user message.
-As shown, the system prompt includes the available tools as a JSON schema.
-Additionally, we also specify the function calling format for using the tools.
-Each call to an LM requires appending an assistant header for completion (not shown here).
-After execution of the tools, we append the tool response, and the assistant continues with the next completion.
-From a probabilistic perspective, given an input $\mathbf{x}=(x_{1},\dots,x_{T})$ (also called a prompt) consisting of a sequence of tokens $x_{t}\in\mathcal{V}$, a post-trained LM defines a distribution over output sequences $\mathbf{y}=(y_{1},\dots,y_{N})$ conditioned on $\mathbf{x}$.
-The LM assistant generates $\mathbf{y}$ in an autoregressive manner:
+\__box_backend_scale:Nnn System You have access to the following functions.
+Respond in JSON format.
+{ "name": "search_web", "description": "Search the web for current information", "parameters": { "type": "object", "properties": { "query": { "type": "string", "description": "The search query" } }, "required": ["query"] } } You are an assistant that provides the user with information from the web.
+User Find the latest trip of IShowSpeed Assistant Tool Response IShowSpeed’s most recent trip was a high-profile tour across China and Mongolia in March and April 2025.
+Figure 4: Example of chat history in LLaMA-3.
+Adapted from [61].
+A chat history starts with a prompt comprising a system and user message in LLaMA-3.
+The system message enumerates available tools (as JSON schemas) and sets high-level behavior.
+The user message specifies the task and the data.
+During the multi-step process, the chat history is updated with assistant messages (LM responses), user feedback, and tool responses.
+From a probabilistic perspective, given an input (corresponding to the chat history) $\mathbf{x}=(x_{1},\dots,x_{T})$ consisting of a sequence of tokens $x_{t}\in\mathcal{V}$, a post-trained LM generates a response $\mathbf{y}=(y_{1},\dots,y_{N})$ conditioned on $\mathbf{x}$ in an autoregressive manner:
 $p_{\theta}(\mathbf{y}\mid\mathbf{x})=\prod_{t=1}^{N}p_{\theta}(y_{t}\mid\mathbf{x},y_{<t}).$
-Post-training defines a broad objective where the LM is trained to be helpful, harmless, and honest Grattafiori et al.
-(2024).
-Typically, we use a mixture of supervised finetuning and preference learning Grattafiori et al.
-(2024).
-Notably, the prompt $\mathbf{x}$ consists of tasks such as function calling and instruction following.
+During post-training, the LM is trained to follow a chat format.
+Thus, the response follows the chat format and corresponds to a tool call or a response to the user.
 As shown in Fig.
-3, the input $\mathbf{x}$ corresponds to a chat where the sources of each message (typically, system, user, assistant, and tool) are specified by special headers.
-At inference, a user provides a query $\mathbf{q}$ and we parse it into a chat $\mathbf{x}$ with a system message and the query, optionally with a pre-appended user input template.
-In principle, at each step, the chat is updated with new messages from the assistant or tool.
-The system message contains, for instance, the general instructions, safety guidelines, constraints, and tools available.
-Additionally, post-trained LMs use function calling to interact with tools in principled ways.
-In formula, we denote a scaffolded LM as $\pi_{\phi}(\mathbf{q})$, where $\phi$ refers to non-parametric variables, including the system prompt and the user input template for parsing $\mathbf{q}$ into $\mathbf{x}$.
+4, the multi-step process in a scaffolded LM produces a chat history where messages come from different roles.
+Specifically, a chat format is a protocol that specifies how to convert a chat history into a single sequence of tokens $\mathbf{x}$ that the post-trained LM both expects as input and emits as output.
+For instance, the “tool” role is used for tool responses.
+In other words, we can understand a chat format as a protocol consisting of
+  • 1.
+    a finite set of roles, e.g., system, user, assistant, tool.
+  • 2.
+    a set of special tokens that act as markers, such as role headers, channel indicators, message delimiters, tool-call fences, and termination markers [62].
+Importantly, the system message and user message specify how the model should behave as well as when and how to use tools.
+A chat starts with a system message and a user message, and we refer to them as the prompt.
+Non-parametric training steers the behavior of scaffolded LMs by optimizing the prompt and tools.
+By default, it is recommended to adhere to the provider’s native chat format with explicit role headers (system, user, assistant, tool), placing tool outputs under the dedicated tool role and assistant generations under the assistant role.
+This aligns with how post-trained LMs are supervised and preserves the intended privilege ordering [63].
+During post-training, the LM is explicitly trained to treat each message source (system → user → tool output) as having an explicit privilege order [63].
+Thus, if there’s a conflict between messages from different sources, the LM is trained to prioritize them based on this privilege order.
+This avoids prompt-injection [64], jailbreaks [65], and system-prompt extraction [66].
+This process is different from knowledge conflict [67, 68] that focuses on factuality.
+It arises when there’s a conflict between parametric knowledge (what the model “knows”), contextual knowledge (retrieved content), and instruction knowledge (user/system assumptions or edits).
+Accordingly, Claude’s system prompt includes explicit instructions to avoid invoking external tools on topics the model already “knows,” which could be a strategy to avoid knowledge conflicts to some extent [66, 69].
 3.2 Tools
-We follow the definition of Wang et al.
-Wang et al.
-(2024e) and define an LM-used tool as a function interface to a computer program that runs externally to the LM, where the LM generates the function calls and input arguments in order to use the tool.
-In that sense, a tool can be virtually anything, including the interface to a digital system, a scratch pad for thinking Anthropic (2025), a request to the user LangChain (2025b), another LM OpenAI (2025), or developer-defined code.
-Notably, the function calling corresponds to a structured output from the post-trained LM that generates function calls and input arguments.
-Function calling provides a principled way to use tools in the sense that it allows the output of the LM to be parsed and provides a unified way to describe available tools to the LM.
-Thus, we follow previous work and use the terms function calling and tool interchangeably Grattafiori et al.
-(2024).
-It is important to note that the order of arguments in tools, as well as the order of tools when calling multiple tools, matter because post-trained LMs are autoregressive models McCoy et al.
-(2023); Grosse et al.
-(2023).
-In practice, a post-trained LM comes with built-in tools or developer-defined code as tools.
-As we will see in Sec.
-3, training of tools refers to updating developer-defined code.
-Built-in tools OpenAI (2025); Anthropic (2025) typically includes:
-  • a web search API Nakano et al. (2022) allows the LM to tackle tasks with public data. The LM typically generates a response by citing the retrieved content.
-  • a vector database Lewis et al.
-(2020); Izacard & Grave (2021); Ram et al.
-(2023); Anthropic (2024); Jiang et al.
-(2020) for storing data as embeddings that can be retrieved through top-$k$ retrieval to a query.
-It is often used for accessing private data or as a memory module Park et al.
-(2023).
-  • a computer use tool OpenAI (2025) that interacts with a headless or headful browser to perform browsing tasks.
-  • a thinking tool Anthropic (2025), unlike ReAct Yao et al.
-(2023b), defines reasoning itself as a callable tool, enabling structured, mid-process reflection that offers a more adaptable alternative—allowing the language model to interleave reasoning steps with multiple tool calls to maintain coherence and improve complex task handling.
+A tool can be virtually anything, including the interface to a digital system, a scratch pad for thinking [70], a request to the user [71], another LM [72], or developer-defined code.
+We define a tool as an external capability that the LM can invoke via the tool API [7, 8].
+Moreover, we define a function as a tool defined by a JSON schema.
+Similarly, we define tool calling as a specific type of response from the LM that includes a call to one of the available tools and a corresponding plain text for that tool (e.g., a query to a search engine).
+We define function calling as a specific type of tool call that corresponds to a structured JSON.
+It is important to note that the order of arguments in a function matters because post-trained LMs are autoregressive models [73, 74].
+During post-training, LMs may be trained to use built-in tools, which we call tool-integrated reasoning [75].
+Built-in tools are executed remotely on the provider side within temporary secure sandboxes, rather than on the user’s local machine.
+Thus, built-in tools are only used for seeking external information to improve their reasoning process rather than advancing on the task itself.
+Built-in tools [23, 70] typically include:
+  • a web search API [76] allows the LM to tackle tasks with public data. The LM typically generates a response by citing the retrieved content.
+  • a vector database [77, 78, 79, 80, 81] for storing data as embeddings that can be retrieved through top-$k$ retrieval for a query.
+It is often used for accessing private data or as a memory module [82].
+  • a computer use tool [83] that interacts with a headful browser. Unlike the web search API, it allows the LM to use interactive websites.
+The multi-step process in scaffolded LMs naturally allows sequential tool composition, where tools are called sequentially in a coordinated manner to solve a task.
+In that sense, scaffolded LMs start by producing a plan in plain text to improve tool composition [75].
+CodeAct [84] proposes to use executable Python code as a unified way to call tools.
+As shown in Fig.
+6, code inherently supports both control flow (e.g., conditionals and loops) and data flow, allowing intermediate results to be stored and reused via variables when coordinating multiple tool calls.
+They [84] find that it achieves comparable or better performance on sequential tool composition.
+However, this approach is vulnerable to executing unsafe code and requires sandboxing.
+Even with sandboxes, strict capability scoping and comprehensive controls are needed: network and file-system egress restrictions, timeouts, resource quotas, import allow-lists, AST/static analysis, and immutable audit trails.
+Building and operating that infrastructure is non-trivial.
+Moreover, code-injection attacks become more dangerous because untrusted tool outputs (e.g., web pages or file contents) can be stitched into code and executed.
+Additionally, agents with access to the internet are exposed to a wide variety of attacks where tools return instructions, when followed by the LM, leads to leaking personal data [64].
+Besides, LMs may also occasionally generate invalid tool calls (e.g., syntax errors), wrong tools, or irrelevant tools.
+For instance, irrelevant tool calls may be a result of reward hacking during post-training [85].
+Similarly, the LM may face unexpected tool execution failures (e.g., rate-limit).
+To account for these, developers must implement error-handling mechanisms to catch exceptions during tool execution and provide feedback to the LM for recovery.
+Additionally, the way tools are defined and described inside the system message affects reliability.
+Clear, specific descriptions and intuitive names help the LM choose the right tool.
+Additionally, clear description provides explicit guidance for the LM to learn when to use this tool as opposed to learning it through RL which is prone to reward hacking [85].
+Thus, the optimization of prompts and tools is important.
 3.3 Multi-Step Process
 Figure 5: Multi-step process of a scaffolded LM.
-The distinctive trait between agent and workflow is that for each workflow step, the developer controls the system prompt and user instruction to pre-define the sub-task at that step.
+The distinctive trait between agent and workflow is that for each workflow step, a different prompt is written to pre-define the sub-task at that step.
 This makes workflows much more efficient and predictable than agents.
@@
 In this figure, we assume the case where the tool execution happens on the developer side.
-For built-in tools, tool execution happens on the model (API) side OpenAI (2025).
-A scaffolded LM involves a multi-step process with tools, where at each step, the LM outputs tool calls.
+For built-in tools, tool execution happens on the provider’s side [23].
+A scaffolded LM involves a multi-step process with tools.
+Each step consists of an interaction between the assistant and tool.
 As shown in Fig.
-4, guiding the model at each step with tailored prompts reduces its decision-making burden.
-Thus, in practice, a workflow involves pre-defining the number of steps and crafting appropriate user or system prompts for each, while an agent focuses on engineering the system prompt and the set of tools.
-These approaches mainly differ in how much autonomy the LM has, and we can also refer to this autonomy as the degree of agenticness LangChain (2025a).
+5, guiding the model at each step with tailored prompts reduces its decision-making burden.
+Thus, in practice, a workflow involves pre-defining the number of steps and crafting appropriate prompt for each step, while an agent focuses on engineering a single prompt and the set of tools.
+These approaches mainly differ in how much autonomy the LM has, and we can also refer to this autonomy as the degree of agenticness [86].
 An agent works well in unstructured scenarios, i.e., in cases where it is hard to anticipate all the scenarios that the LM might face.
 In contrast, a workflow approach works well when we know the solution path to a problem and can thus directly define each step required to solve it.
-A good example is solving GitHub issues in SWE-Bench Jimenez et al.
-(2024).
-A workflow such as Agentless Xia et al.
-(2024) solves GitHub issues with the following steps.
+A good example is solving GitHub issues in SWE-Bench [50].
+A workflow such as Agentless [49] solves GitHub issues with the following steps.
 Roughly, it first calls the LLM with the GitHub repository and issue description to provide a list of files that potentially contain the bug.
@@
 Then, it asks the LLM to generate multiple patches.
-In comparison, an agent such as Openhands Wang et al.
-(2024c) provides the LM with a terminal and a code interpreter to interact with the repository.
+In comparison, an agent such as Openhands [5] provides the LM with a terminal and a code interpreter to interact with the repository.
 3.3.1 Agents
@@
 They also determine when a task is solved based on their own interpretation of results and goals.
-An agent starts with a system message and a user message and generates the first response.
-Then, at each step, it appends to the input $\mathbf{x}$ the LM response and tool execution response, and generates the next response provided with the tool execution response.
-It continues until it decides to stop or it reaches the maximum number of steps, and (optionally) returns a response to the user.
-Note that while Fig.
-3 appends the LM response and tool response under the special headers, we can also append them to the user message.
-Besides providing the right tool to our agent, an important design of the agent is how it generates function calls and input arguments.
-In particular, it affects the horizon as well as the task complexity.
-Traditionally, people have used structured formats such as JSON or XML to generate tool calls.
-Recently, CodeAct Wang et al.
-(2024b) proposes to use executable Python code as a unified way to call tools.
+The chat of an agent starts with a prompt and each step updates the chat history.
+The agent continues until it decides to stop or it reaches the maximum number of steps, and returns a response to the user that summarize what it did.
+Besides providing the right tool to our agent, an important design of the agent is how it composes tools.
+In particular, code execution [84] offers a unified way to compose tools with control flow and data flow, and typically results in a shorter horizon.
+Another approach prompts the agent to provide a plan beforehand [75].
+3.3.2 Workflows
+Workflows [49, 87, 88] shine in applications where a solution path can be pre-determined.
+In contrast to agents, they are rendered more predictable and reliable by limiting the LM’s autonomy.
+At each step, the input $\mathbf{x}$ consists of a different prompt that specify the sub-task at that step.
+Thus, workflows consist of a pre-determined number of steps.
+It is more cost-efficient as it avoids any unnecessary exploration steps [49].
+Unlike agents that scale with the number of steps [89], workflows scale by generating multiple responses in parallel [90, 56, 91].
+Specifically, by fixing the solution path, the output format at each step is predictable.
+Thus, workflows use meta-generation techniques [56] such as best of $n$ sampling or majority voting [92] to scale test-time compute.
+For instance, in [88], the AI scientist-v2 workflow can reach workshop-level papers by implementing tree-based methods on top of its previous version [87].
+In particular, by generating multiple solutions at each step, it effectively generates multiple solution paths, and then uses heuristics (e.g., LLM as judge [93]) to select the best-performing path.
+The cost of parallelization in workflows mainly depends on the aggregation method.
+For instance LLM as judge adds one extra LM call, while pairwise judging adds quadratic cost.
+Meanwhile, aggregation can be done without an LM judge, e.g., unit tests [49] or majority voting [92].
+Figure 6: Tool compositionality.
+Left: tools are composed sequentially.
+As shown in the figure, some LMs support generating independent tool calls in parallel [7].
+Right: executable Python code as a unified way to call tools [84].
+Code inherently supports both control flow (e.g., conditionals and loops) and data flow, allowing intermediate results to be stored and reused via variables when coordinating multiple tool calls
+3.4 Benchmarks
+We highlight a growing trend in recent benchmarks: from measuring pure intelligence (e.g., math olympiads) to measuring productivity gains (e.g., software engineering) [4].
+The goal is no longer to assess whether a model solves a task perfectly in isolation, but how reliably and usefully it supports human workflows.
+Importantly, these benchmarks are used to compare LMs on products [94, 5, 95, 96] that operate in mixed-autonomy.
+This is a setting where humans and AI work together, sharing control and decision-making responsibilities to achieve a goal.
+For instance, in vibe coding, the human guides intent and judges quality, while the AI proposes implementations and refactors.
+To accurately gauge mixed-autonomy capability, these benchmarks align agent performance with human effort—measuring monetary return [97], simulating human collaborators [98, 99], exercising real software and APIs [50, 100, 101], and using time as a measure of task difficulty [102] because time is money.
+SWE-bench [50] and SWE-Lancer [97] ground evaluation in software engineering tasks from GitHub or Upwork—tasks that are compensated in real-world settings and evaluated via functional test cases.
+Other benchmarks go beyond measuring raw accuracy.
+BrowseComp [100] evaluates accuracy and calibration error.
+In particular, the LM is asked to output a confidence (0–100) with its answer; BrowseComp then evaluates how well that confidence aligns with reality.
+HCAST [102] evaluates success rate as a function of human-calibrated task length (e.g., <15min, 15min-15h, 1-4h, 4h+).
+Additionally, they compute a “50% time-horizon”: the human task duration at which the agent has a 50% chance of success, fitted from per-task success vs.
+human time.
+RE-Bench [103] finds that currently humans scale better with increased time budgets to solve the task.
+For example, $\tau$-Bench [98] and Sweet-RL [99] simulate human feedback using LMs to better reflect the back-and-forth nature of co-working.
+For tasks that are hard to verify, PaperBench [101] proposes a rubric organized as a hierarchical tree of requirements.
+The leaf nodes are precise, binary checks (pass/fail).
+Internal nodes aggregate their children by a manual, importance-based weighting, and the root score is the paper’s Replication Score.
+Tab.
+1 summarizes these benchmarks.
+Benchmark Scenario Evaluation metrics SWE-Bench [50] Software engineering Unit test.
+SWE-Lancer [97] Software engineering pass@1 / pass@k, dollars earned, unit test, SWE Manager by matching the original manager’s proposal.
+HCAST [102] Machine learning, software engineering, cybersecurity Binarized success rate by human-time buckets.
+RE-Bench [103] AI Research Expert-graded task success; fraction of human pace; time-to-solution.
+PaperBench [101] AI research Rubric organized as a hierarchical tree of requirements.
+$\tau$-Bench [98] Tool-agent-user interaction pass@k metric as a measure of reliability of an agent.
+SWEET-RL [99] Multi-turn collaborative reasoning Task success/win rate under simulated feedback.
+BrowseComp [100] Deep research Accuracy, calibration error.
+Table 1: Benchmarks commonly used for scaffolded LMs.
+4 Training with Language Supervision
+We view scaffolded LM $\pi$ as semi-parametric models where the LM is the parametric component that is fixed.
+Instead, we consider the training of non-parametric variables, prompt and tools, that are represented in language and code.
 As shown in Fig.
-5, compared to structured formats like XML, code inherently supports both control flow (e.g., conditionals and loops) and data flow, allowing intermediate results to be stored and reused via variables when coordinating multiple tool calls.
-They Wang et al.
-(2024b) find that it achieves comparable or better performance on atomic tool use, while excelling in tasks requiring multiple calls to multiple tools in multi-turn interactions.
-However, it requires a sandboxed environment for code execution.
-2.3.2 Workflows
-Workflows Xia et al.
-(2024); Lu et al.
-(2024); Yamada et al.
-(2025) shine in applications where a solution path can be pre-determined.
-In contrast to agents, they are rendered more predictable and reliable by limiting the LM’s autonomy.
-At each step, the input $\mathbf{x}$ consists of a system message and a user message that is updated to specify the sub-task at each step.
-Thus, workflows consist of a pre-determined number of steps.
-It is more cost-efficient as it avoids any unnecessary exploration steps Xia et al.
-(2024).
-Unlike agents that scale with the number of steps Pan et al.
-(2024), workflows scale by generating multiple responses in parallel Snell et al.
-(2024); Welleck et al.
-(2024); Chen et al.
-(2024a).
-Specifically, by fixing the solution path, the output format at each step is predictable.
-Thus, workflows use meta-generation techniques Welleck et al.
-(2024) such as best of $n$ sampling or majority voting Wang et al.
-(2023b) to scale test-time compute.
-For instance, in Yamada et al.
-(2025), the AI scientist-v2 workflow can reach workshop-level papers by implementing tree-based methods on top of its previous version Lu et al.
-(2024).
-In particular, by generating multiple solutions at each step, it effectively generates multiple solution paths, and then uses heuristics (e.g., LLM as judge Yao et al.
-(2023a)) to select the best-performing path.
-Figure 5: Generating function calls and input arguments.
-We show an agent that uses a tool with structured format or executable code Wang et al.
-(2024b).
-Provided with a system prompt to control how it interacts with the tool, the model either generates output that is easily parsed or directly generates Python code that executes the tool.
-2.4 Benchmarks
-As scaffolded language models (LMs) become increasingly integrated into real-world Windsurf ; Wang et al.
-(2024c); Cursor (2023); Microsoft (2025), there is growing momentum toward developing benchmarks that reflect mixed autonomy settings—scenarios where humans and AI share control and collaborate to achieve shared goals.
-In practice, these models are not isolated problem-solvers; they operate through the same APIs, UIs, and tools as human experts, and seldom succeed on the first try Rein et al.
-(2025); Wijk et al.
-(2024); Wei et al.
-(2025a); Yoran et al.
-(2024); Jimenez et al.
-(2024); Pan et al.
-(2024); Zhong et al.
-(2025); Jha et al.
-(2025); Starace et al.
-(2025).
-This necessitates multi-turn interactions, human-in-the-loop feedback, and iterative refinement—dynamics absent from traditional benchmarks focused on raw intelligence (e.g., math Olympiads or code generation).
-To evaluate real-world effectiveness, recent benchmarks have shifted from measuring standalone task accuracy to assessing human-AI collaboration and productivity gains.
-For example, Tau-Bench Yao et al.
-(2024a) and Sweet-RL Zhou et al.
-(2025) simulate human feedback using LMs to better reflect the back-and-forth nature of co-working.
-Benchmarks like HCAST Rein et al.
-(2025) and RE-Bench Wijk et al.
-(2024) calibrate task difficulty against human effort—reporting metrics like “fraction of human pace” rather than raw accuracy.
-SWE-Bench Jimenez et al.
-(2024) and SWE-Lancer Miserendino et al.
-(2025) go further, grounding evaluation in software engineering tasks from GitHub or Upwork—tasks that are compensated in real-world settings and evaluated via functional test cases.
-This benchmarking evolution reflects a shift: from measuring intelligence to measuring impact OpenAI (2018).
-The goal is no longer to assess whether a model solves a task perfectly in isolation, but how reliably and usefully it supports human workflows.
-Mixed autonomy systems must fluidly ratchet delegation up or down, adapt to under-specified tasks, navigate unclear success criteria, and coordinate across asynchronous streams of work.
-As highlighted by the METR report, such messiness is correlated with task length and reflects the realities of intellectual labor.
-Accordingly, we argue for new evaluation paradigms that go beyond single-turn metrics and instead quantify collaborative efficiency, reliability, and human-labor savings in dynamic, real-world environments.
-3 Training with Language Supervision
-We view scaffolded LM as semi-parametric models where the LM is the parametric component $\theta$ that is fixed.
-Instead, we consider the training of non-parametric variables $\phi$ that are represented in language or code.
-Thus, we denote our scaffolded LM as $\pi_{\phi}$, where we omit $\theta$ as it is fixed.
+8, non-parametric variables include the prompt and developer-defined code as tools.
 As shown in Fig.
-7, non-parametric variables $\phi$ include the system prompt, the user input template, developer-defined code as tools, or the scaffold’s code.
-As shown in Fig.
-6, we use an LM as an optimizer to generate the updated variables $\phi$ according to an objective $J(\phi)$, an execution trace $\tau$, and optionally execution feedback $f$, all provided in natural language.
+7, we use an LM as an optimizer to generate the updated prompt or tool according to an objective (“improve prompt according to feedback on this execution trace”), an execution trace, and optionally execution feedback $f$, all provided in natural language.
 As in machine learning (ML), we assume a training, validation, and test dataset sampled from the same (unknown) distribution.
-During training, we iteratively sample data point(s) and update $\phi$.
-At inference, $\phi$ is fixed.
-Unlike in ML, where we train with Monte Carlo samples to prevent catastrophic forgetting Hadsell et al.
-(2020), an LM optimizer works with interpretable variables $\phi$ where the role of batching is still an open research question Yuksekgonul et al.
-(2025).
-The datasets $\mathcal{D}$ consist of natural language queries $\{\mathbf{q}\}$ corresponding to user input from which we obtain execution traces $\tau\sim\pi_{\phi}(\mathbf{q})$.
-During training, the execution traces come from any scaffolded LM $\pi_{\phi^{\prime}}$ or the current scaffolded LM $\pi_{\phi}$.
-In that sense, it shares similarity with RL where we have on/off-policy training, offline/online training Sutton & Barto (2018).
+During training, we iteratively sample data point(s) and update prompt or tool.
+At inference, prompt(s) and tools are fixed (in the case of workflows there’s a multiple prompts, i.e., a different prompt for each step).
+Unlike in ML, where we train with Monte Carlo samples to prevent catastrophic forgetting [3], an LM optimizer works with variables represented in language or code as opposed to tensors.
+Thus, the role of batching is still an open research question [9].
+The datasets $\mathcal{D}$ consist of user message corresponding to different tasks from which we obtain execution traces $\tau\sim\pi$.
+During training, the execution traces come from any scaffolded LM $\pi^{\prime}$ or the current scaffolded LM $\pi$.
+In that sense, it shares similarities with RL, where we have on/off-policy training and offline/online training [104].
 However, the training dynamics significantly differ from those in RL.
-While RL is about finding good data (exploration) and imitating that good data (exploitation) Eysenbach et al.
-(2020), recent works in Sec.
+While RL is about finding good data (exploration) and imitating that good data (exploitation) [105], recent works in Sec.
 4.3 learn only from bad execution traces.
@@
 Optionally, execution feedback can be provided.
-Central to this paradigm is an optimizer implemented with an LM as opposed to bootstrapping successful execution traces as in-context learning examples Khattab et al.
-(2022); Von Oswald et al.
-(2023).
-We further discuss how this paradigm differs from other approaches that have employed an LM and non-parametric components for improving a scaffold LM.
-First, in Reflexion Shinn et al.
-(2023), the updates are not retrained across tasks and are only used for retrying a task.
-Second, CoALA Sumers et al.
-(2024) is an orthogonal approach that works at run-time with a specific scaffold.
+Central to this paradigm is an optimizer implemented with an LM as opposed to bootstrapping successful execution traces as in-context learning examples [106, 107].
+We further discuss how this paradigm differs from other approaches that have employed an LM and non-parametric components for improving a scaffolded LM.
+First, in Reflexion [108], the updates are not retrained across tasks and are only used for retrying a task.
+Second, CoALA [109] is an orthogonal approach that works at run-time with a specific scaffold.
 In particular, it requires changing the design of the agent with a memory module.
-This memory module not only serves as long-context but also for improving the agent over time.
+This memory module not only serves as long-context memory but also for improving the agent over time.
 However, this requires the agent to make additional decisions on what and when to store or retrieve knowledge from this memory module.
 In training with language supervision, recent works focus on a setting where we are provided with a fixed dataset for training.
-Thus, it can be seen as amortizing the expensive inference of cognitive architecture Sumers et al.
-(2024) into a training procedure.
+Thus, it can be seen as amortizing the expensive inference of a cognitive architecture [109] into a training procedure.
 Our organization follows the structure outlined in Fig.
-7.
+8.
 Tracing their development, these methods have evolved to optimize more complex execution traces.
@@
 4.1 Prompt Optimization
-In prompt optimization Li et al.
-(2025b), we update the instruction $\phi$ of an LM (as opposed to a scaffolded LM).
-We prompt an LM optimizer with an objective $J(\phi)$ and (optionally) a textual feedback $f$.
-In its simplest form, we provide query and response examples and prompt the LM to generate the corresponding instruction Khattab et al. (2023); Zhou et al. (2023).
+In prompt optimization [110], we update the prompt of an LM (as opposed to a scaffolded LM).
+We prompt an LM optimizer with an objective (e.g., “improve this instruction according to the following examples of tasks where the model fails”) and (optionally) a textual feedback $f$.
+The updates made by the LM optimizer is evaluated on a validation set and the best one is selected.
+Thus, while these methods use a different objective in natural language, by heuristically selecting the best update on a validation set, they implicitly optimize for performance on the validation set.
+In its simplest form, we provide query and response examples and prompt the LM to generate the corresponding instruction [20, 21].
 Provided with different initial instructions, other methods define the objective as generating variations.
 Each variation is evaluated on a validation set, and the top-performing versions are retained.
-In Zhou et al.
-(2023), the objective is to generate semantically similar instructions.
-Evolutionary methods Fernando et al.
-(2024); Guo et al.
-(2024) combine parts of multiple instructions and then introduce random changes; we call this process crossover and mutation.
-In OPRO Yang et al.
-(2024), we define the objective as generating instructions with a higher score using previous instructions, their scores, and a random sample of input-output pairs from the training data.
-To introduce fine-grained variations of instructions, SAMMO Schnabel & Neville (2024) uses a more granular representation of prompts by breaking them down into smaller sections, such as input formatting.
+In [21], the objective is to generate semantically similar instructions.
+Evolutionary methods [111, 112] combine parts of multiple instructions and then introduce random changes; we call this process crossover and mutation.
+In OPRO [113], we define the objective as generating instructions with a higher score using previous instructions, their scores, and a random sample of input-output pairs from the training data.
+To introduce fine-grained variations of instructions, SAMMO [114] uses a more granular representation of prompts by breaking them down into smaller sections, such as input formatting.
 This fine-grained structure allows SAMMO to adjust specific components, such as changing the input format from raw text to a structured format (e.g., XML).
-In a similar vain, ABO Ma et al.
-(2024) implements the instruction as step-by-step guidelines.
+In a similar vein, ABO [115] implements the instruction as step-by-step guidelines.
 Thus, the LM optimizer only updates the specific step that failed.
@@
 LM optimization approaches have evolved from prompt optimization, then agents, and finally workflows.
-In all approaches, an LM acts as an optimizer for generating updates on $\phi$.
+In all approaches, an LM acts as an optimizer for generating updates on prompts or tools.
 Top: prompt optimization focuses on LMs as opposed to scaffolded LMs.
 The LM optimizer uses a single or multiple LM calls.
-Middle Experiential Learning extends to agents and focuses on extracting insight from tasks and refining tools.
-The LM optimizer uses a single LM call to update them from successful/unsuccessful execution traces.
-Bottom: AutoDiff Frameworks focus on workflows where computation is represented as a graph.
+Middle: experiential learning extends to agents and focuses on extracting insight from tasks and storing them in the system message, or refining tools.
+The LM optimizer uses a single LM call to update them from successful or unsuccessful execution traces.
+Bottom: AutoDiff frameworks focus on workflows where computation is represented as a graph interleaving LM calls and developer-defined code as tools.
 Using textual feedback, the LM optimizer can optimize nodes of the graph (e.g., system prompts).
-Another line of work uses an LM to generate rich textual feedback $f$, and defines the objective as following the feedback $f$ Pryzant et al.
-(2023); Wang et al.
-(2024d); Chen et al.
-(2024c); Austin & Chartock (2024).
+Another line of work uses an LM to generate rich textual feedback $f$, and defines the objective as following the feedback $f$ [39, 38, 116, 117].
 Importantly, textual feedback is obtained from sample queries where the target LM fails.
-Additionally, these works also propose summarizing multiple pieces of feedback Chen et al.
-(2024c) or updates Austin & Chartock (2024).
-Self-discover Zhou et al.
-(2024a) focuses on the output structure of the instruction.
+Additionally, these works also propose summarizing multiple pieces of feedback [116] or updates [117].
+Self-discover [118] focuses on the output structure of the instruction.
 The objective is to generate a JSON format with keys representing distinct reasoning steps.
@@
 4.2 Experiential Learning
-We borrow the term experiential learning from ExpeL Zhao et al.
-(2024a) to refer to the training of agents with language supervision.
-The LM optimizer considers an objective $J(\phi)$ (e.g., “improve set of tools according to feedback”, “extract reusable tool from execution trace”), user queries $\mathbf{q}$, execution trace $\tau$, optionally a textual feedback $f$, and generates the updated variable $\phi^{\prime}$.
-In particular, the execution trace $\tau$ of agents consists of a sequence of tool calls/responses and may end with a response to the user.
+We borrow the term experiential learning from ExpeL [10] to refer to the training of agents with language supervision.
+The LM optimizer considers an objective (e.g., “extract a general insight on why you failed this task”, “extract reusable tool from execution trace”), prompt $\mathbf{s}_{0}$, execution trace $\tau$, optionally a textual feedback $f$, and generates an insight or a tool.
+In particular, the execution trace $\tau$ of agents consists of a sequence of tool calls/responses and ends with a response to the user.
 As discussed in Sec.
-2.3.1, tools play an important role in agents.
+3.3.1, tools play an important role in agents.
 As shown in Fig.
-8, $\phi$ is either the set of tools or insights that are provided to the agent’s system prompt.
+9, the objective corresponds to either improving the set of tools or extract insights that are provided to the agent’s system prompt.
 Specifically, prior works leverage successful and unsuccessful execution traces with different objectives.
 4.2.1 Insights Extraction
-ExpeL Zhao et al.
-(2024a) starts by collecting a dataset for training consisting of user queries $\mathbf{q}$ and corresponding execution trace $\tau$.
-Using a Reflexion agent Shinn et al.
-(2023), multiple execution traces may be sampled for a query $\mathbf{q}$.
-In particular, the Reflexion agent allows the agent to retry the task upon failing, which allows the collection of unsuccessful and successful trajectories $(\tau^{\text{succ}},\tau^{\text{fail}})$ on the same task instance $\mathbf{q}$.
-The LM optimizer refines a list of insights from successful and unsuccessful execution traces $(\tau^{\text{succ}},\tau^{\text{fail}})$ for a given query $\mathbf{q}$ or from $k$ successful trajectories $(\tau_{1}^{\text{succ}},...,\tau_{k}^{\text{succ}})$.
+ExpeL [10] starts by collecting a dataset for training consisting of prompt $\mathbf{s}_{0}$ and corresponding execution trace $\tau$.
+Using a Reflexion agent [108], multiple execution traces may be sampled for a prompt $\mathbf{s}_{0}$.
+In particular, the Reflexion agent allows the agent to retry the task upon failing, which allows the collection of unsuccessful and successful trajectories $(\tau^{\text{succ}},\tau^{\text{fail}})$ on the same prompt instance $\mathbf{s}_{0}$.
+The LM optimizer refines a list of insights from successful and unsuccessful execution traces $(\tau^{\text{succ}},\tau^{\text{fail}})$ for a given prompt $\mathbf{s}_{0}$ or from $k$ successful trajectories $(\tau_{1}^{\text{succ}},...,\tau_{k}^{\text{succ}})$.
 At inference, those insights are fixed and provided to the scaffolded LM through its system prompt.
-Follow-up works on ExpeL Zhao et al.
-(2024a), such as AutoGuide Fu et al.
-(2024b), modify insights to be selectively added to the context at each step.
+Follow-up works on ExpeL [10], such as AutoGuide [119], modify insights to be selectively added to the context at each step.
 Insights are specific to a state and contain a description of the state in which it is applicable.
 However, these methods require additional cost at inference since an LM summarizes the current state and then uses it to select the corresponding state-specific insight.
-AutoManual Chen et al.
-(2024b) modifies ExpeL Zhao et al.
-(2024a) to prevent distributional shift in offline RL, where at test-time the state distribution is different Wang et al.
-(2023a).
-However, their paper does not show strong evidence that training agents from language supervision suffers from distributional shift Wang et al.
-(2023a).
-Additionally, MSI Fu et al.
-(2024a) proposes to categorize each insight and retrieve the relevant one at test-time based on pre-defined rules.
-ExpeL Zhao et al.
-(2024a) also explores new settings such as transfer learning, where we use insights learned from a source dataset $\{\mathbf{q}^{\text{source}}\}$ and transfer it to a target dataset $\{\mathbf{q}^{\text{target}}\}$.
-In parametric training Pouyanfar et al.
-(2018), it consists of initializing the weights with the model trained on the source dataset.
+AutoManual [120] modifies ExpeL [10] to prevent distributional shift in offline RL, where at test-time the state distribution is different [121].
+However, their paper does not show strong evidence that training agents from language supervision suffers from distributional shift [121].
+Additionally, MSI [122] proposes to categorize each insight and retrieve the relevant one at test-time based on pre-defined rules.
+ExpeL [10] also explores new settings such as transfer learning, where we use insights learned from a source dataset $\{\mathbf{s}_{0}^{\text{source}}\}$ and transfer it to a target dataset $\{\mathbf{s}_{0}^{\text{target}}\}$.
+In parametric training [123], it consists of initializing the weights with the model trained on the source dataset.
 Then, finetuning these weights on the target dataset.
@@
 Additionally, they manually inspect execution traces from the agent with and without the insights.
-For instance, in ALFWorld Shridhar et al.
-(2020), one insight would be “when searching for an item, consider its nature and its typical usage.” This leads the agent to update its belief Andreas (2022) and avoid unnecessary actions to find objects.
+For instance, in ALFWorld [124], one insight would be “when searching for an item, consider its nature and its typical usage.” This leads the agent to update its belief [125] and avoid unnecessary actions to find objects.
 Furthermore, they find that another insight, “if an attempt to interact with an item fails […] consider alternative actions or locations,” leads to self-correction.
@@
 Figure 9: Experiential learning focuses on tools and insights.
-Successful execution traces are used for extracting reusable tools Wang et al.
-(2025c) while unsuccessful execution traces are used to refine tools Zhang et al.
-(2024).
-Additionally, we can contrast successful and unsuccessful trajectories on the same query to extract insights Zhao et al.
-(2024a).
+Successful execution traces are used for extracting reusable tools [11] while unsuccessful execution traces are used to refine tools [126].
+Additionally, we can contrast successful and unsuccessful trajectories on the same task to extract insights [10].
 4.2.2 Tools Optimization
-Tool optimization Zhang et al.
-(2024); Zhao et al.
-(2024c); Wang et al.
-(2025c) focuses on refining developer-defined code into more effective routines.
-The main challenge of tool optimization is ensuring that the skills are reusable and useful across episodes Wang et al.
-(2024a).
-A related term to tool, but slightly more general is skill Wang et al.
-(2024f).
+Tool optimization [126, 127, 11] focuses on refining developer-defined code into more effective routines.
+The main challenge of tool optimization is ensuring that the skills are reusable and useful across episodes [128].
+A related term to tool, but slightly more general is skill [129].
 It refers to both a textual representation of a sequence of action and developer-defined code as tool.
@@
 First, we want to make tools more efficient, thereby achieving tasks with fewer steps.
-Second, similarly to unsupervised RL (URL) Zhao et al.
-(2022); Eysenbach et al.
-(2019), we want to extract the set of behaviors possible in the environment.
+Second, similarly to unsupervised RL (URL) [130, 131], we want to extract the set of behaviors possible in the environment.
 Drawing an analogy with principal component analysis, before we are given any tasks, we want to know what is the basis of behavior that exists in some environment.
-Similarly, these methods Liu et al.
-(2025b); Wang et al.
-(2024f) stores skills into a vector database and retrieve corresponding skill at inference.
-For instance, in CER Liu et al.
-(2025b) skills are represented as a sequence of tool calls with a high-level description of the skill.
-In AWM Wang et al.
-(2024f), they explore different textual representation of skills.
-We can also understand these methods as ‘curating’ in-context samples for retrieval of similar experiences in ExpeL Zhao et al.
-(2024a).
-However, the textual representation represents a challenge for formal verification Wang et al.
-(2025c).
-At each step, LearnAct Zhao et al.
-(2024c) and AgentOptimizer Zhang et al.
-(2024) sample an execution trace $\tau\sim\pi_{\phi}(\mathbf{q})$, and refine the set of tools for unsuccessful execution traces $\tau^{\text{fail}}$ LearnAct uses an execution feedback $f$ corresponding to execution errors.
+Similarly, these methods [132, 129] stores skills into a vector database and retrieve corresponding skill at inference.
+For instance, in CER [132] skills are represented as a sequence of tool calls with a high-level description of the skill.
+In AWM [129], they explore different textual representation of skills.
+We can also understand these methods as ‘curating’ in-context samples for retrieval of similar experiences in ExpeL [10].
+However, the textual representation represents a challenge for formal verification [11].
+At each step, LearnAct [127] and AgentOptimizer [126] sample an execution trace $\tau\sim\pi(\mathbf{s}_{0})$, and refine the set of tools for unsuccessful execution traces $\tau^{\text{fail}}$.
+LearnAct uses execution feedback $f$ corresponding to execution errors.
 The LM optimizer’s objective is to address $f$ by either revising, adding a tool, or updating the system prompt with information about the tools.
@@
 However, this is expensive and quickly leads to overfitting, where tools are not reusable.
-In contrast, ASI Wang et al.
-(2025c) extracts reusable tools from a successful execution trace.
+In contrast, ASI [11] extracts reusable tools from a successful execution trace.
 One important challenge in extracting a reusable tool is to make sure that it is both useful and bug-free.
-AgentOptimizer Zhang et al.
-(2024) checks the performance on the training set with and without the updated tools by simulating again on the same query.
+AgentOptimizer [126] checks the performance on the training set with and without the updated tools by simulating again on the same query.
 An LM is used to rewrite the execution trace with only the extracted tools.
 Then, the agent attempts to finish the execution.
-If it successfully solves the task, then the extract tool is kept.
-In contrast to transfer learning for insights, ASI explores domain adaptation Ovadia et al.
-(2019) in web browsing.
+If it successfully solves the task, then the extracted tool is kept.
+In contrast to transfer learning for insights, ASI explores domain adaptation [133] in web browsing.
 They find that the agent can effectively reuse tools and adapt tools to the new domain.
-Similar to unsupervised RL Zhao et al.
-(2022), recent works propose to use an LM to both propose tasks and extract tools from rollouts on these tasks.
-In Voyager Wang et al.
-(2024a), the LM optimizer refines tools based on execution feedback and adds them to the skill set when successfully used to achieve a task.
-SkillWeaver Zheng et al.
-(2025) extracts tool directly from successful execution traces.
+Similar to unsupervised RL [130], recent works propose to use an LM to both propose tasks and extract tools from rollouts on these tasks.
+In Voyager [128], the LM optimizer refines tools based on execution feedback and adds them to the skill set when successfully used to achieve a task.
+SkillWeaver [134] extracts tools directly from successful execution traces.
 Then, it produces test cases to check the long-term viability of the tool.
 4.3 AutoDiff Frameworks
-The term autodiff is inspired by deep learning frameworks Paszke et al.
-(2019) in the sense that it implements elementary operations for defining and optimizing workflows Yuksekgonul et al.
-(2025); Cheng et al.
-(2024); Zhou et al.
-(2024b); Yin (2024).
+The term autodiff is inspired by deep learning frameworks [135] in the sense that it implements elementary operations for defining and optimizing workflows [9, 27, 136, 28].
 The frameworks represent execution traces with a directed acyclic graph (DAG) where nodes correspond to variables or operations and edges represent dependencies.
-They roll out execution traces $\tau\sim\pi_{\phi}$ and obtain execution feedback $f$.
-Subsequently, prior works Yuksekgonul et al.
-(2025); Yin (2024); Zhou et al.
-(2024b); Wang et al.
-(2025a) use the chain rule to back-propagate $f$ through the graph.
+They roll out execution traces $\tau\sim\pi$ and obtain execution feedback $f$.
+Subsequently, prior works [9, 28, 136, 137] use the chain rule to back-propagate $f$ through the graph.
 Each variable of the graph is updated by the LM optimizer according to its feedback.
-Alternatively, Trace Cheng et al.
-(2024) converts the DAG into text akin to a Python traceback.
+Alternatively, Trace [27] converts the DAG into text akin to a Python traceback.
 The LM optimizer updates variables in an autoregressive manner.
@@
 Variables correspond to non-parametric components of the workflow (e.g., system prompt) or intermediate values (result from operations).
-Operations correspond to computation done on variables, e.g., Python function, LM call.
+Operations correspond to computation done on variables, e.g., Python functions, LM calls.
 As an analogy with graph representation for neural networks, variables would be weights, and operations would be layers.
-However, different from neural networks, any node of the computational graph can be optimized, including operations Cheng et al.
-(2024).
+However, different from neural networks, any node of the computational graph can be optimized, including operations [27].
 We distinguish between two lines of work, focusing on prompts or Python functions.
@@
 Thus, we do not consider applications of these frameworks for optimization instances (as opposed to workflows).
-TextGrad Yuksekgonul et al.
-(2025) and other variants Yin (2024); Wang et al.
-(2025a); Zhou et al.
-(2024b) focus on textual variables.
-In particular, operations include concatenations and LM calls Yuksekgonul et al.
-(2025).
+TextGrad [9] and other variants [28, 137, 136] focus on textual variables.
+In particular, operations include concatenations and LM calls [9].
 Accordingly, trainable variables correspond to the system prompt or user input template of an LM call.
-Trace Cheng et al.
-(2024) focuses on general-purpose workflows.
+Trace [27] focuses on general-purpose workflows.
 Importantly, operations are not limited to LM calls, they can be arbitrary Python functions.
@@
 Additionally, they support the optimization of nodes corresponding to operations.
-In that sense, it goes beyond prompts to optimizing the workflow’s code.
+In that sense, it goes beyond prompts to optimizing the workflow’s developer-defined tool.
 Chain Rule
@@
 formatting of execution trace.
-AutoDiff frameworks consist of three key steps: (1) forward computation, (2) backward feedback propagation, (3) variables update.
+AutoDiff frameworks consist of three key steps: (1) forward computation, (2) backward feedback propagation, (3) variable updates.
 Both approaches differ during their backward pass.
@@
 4.3.2 Graph Execution and Optimization
-User input Template $\phi$Input data $\mathbf{q}$concat($\phi$, $\mathbf{q}$)inputLM callresponseevaluate(output)feedback
+system messageuser messageconcat(system message, user message)inputLM callresponseevaluate(output)feedback
 Figure 11: Chain rule in TextGrad.
@@
 It computes a gradient for each variable (in green), provided with gradient functions attached to operations (in yellow).
-We refer to graph execution and optimization as the process of sampling an execution trace $\tau\sim\pi_{\phi}(\mathbf{q})$ and optimizing variables $\phi$ of the graph according to the execution feedback $f$.
+We refer to graph execution and optimization as the process of sampling an execution trace $\tau\sim\pi(\mathbf{s}_{0})$ and optimizing variables of the graph according to the execution feedback $f$.
 As shown in Fig.
-9, it involves three key steps: (1) forward computation, (2) backward feedback propagation, (3) variables update.
-During forward computation, we collect an execution feedback $f$.
-Then, backward feedback propagation performs a reverse topological sort to either apply the chain rule on the execution feedback Yuksekgonul et al.
-(2025); Yin (2024); Wang et al.
-(2025a); Zhou et al.
-(2024b) or format a prompt template with the DAG Cheng et al.
-(2024).
+10, it involves three key steps: (1) forward computation, (2) backward feedback propagation, (3) variables update.
+During forward computation, we collect execution feedback $f$.
+Then, backward feedback propagation performs a reverse topological sort to either apply the chain rule on the execution feedback [9, 28, 137, 136] or format a prompt template with the DAG [27].
 On the one hand, the chain rule enables structured credit assignment by propagating feedback through the graph in a fine-grained manner.
@@
 As shown in Fig.
-10 and following the graph abstraction presented in Sec.
-3.3.1.
-Operations, including LM call and concatenation, create nodes and attach a grad function.
+11, and following the graph abstraction presented in Sec.
+4.3.1, operations, including LM calls and concatenation, create nodes and attach a grad function.
 Most importantly, the grad function contains a prompt for an LM that informs about that operation’s role.
@@
 For each trainable variable, the LM optimizer’s objective follows the feedback as well as the variable’s role.
-Follow-up work extends Textgrad Yuksekgonul et al.
-(2025) in the following ways.
-AdalFlow Yin (2024) extends operations to retrievers Lewis et al.
-(2020).
-GASO Wang et al.
-(2025a) modifies the gradient operation to take into account neighborhood nodes.
-A conceptually more efficient approach is Trace Cheng et al.
-(2024).
+Follow-up work extends Textgrad [9] in the following ways.
+AdalFlow [28] extends operations to retrievers [77].
+GASO [137] modifies the gradient operation to take into account neighboring nodes.
+A conceptually more efficient approach is Trace [27].
 At each step of the optimization process, the chain rule requires multiple calls to an LM.
 Furthermore, the LM optimizer also requires multiple calls to an LM.
-In contrast, Trace Cheng et al.
-(2024) formats a prompt template for the execution trace during backward feedback propagation.
+In contrast, Trace [27] formats a prompt template for the execution trace during backward feedback propagation.
 This means the LM optimizer generates all updated variables in an autoregressive manner, provided with the execution trace and feedback.
 Additionally, as discussed in Sec.
-3.3.1, Trace Cheng et al.
-(2024) uses a different graph abstraction that allows for arbitrary Python functions as operations.
-As opposed to Textgrad Yuksekgonul et al.
-(2025), this allows the LM optimizer to also consider operations during training.
-In that sense, we can optimize the workflow’s code.
-4 Discussion
-In the context of rapid progress in the field of LMs, it is important to ground our discussion to the broader research in LMs, specifically on the recent breakthroughs in LMs focusing on training paradigms that elicit general capabilities Sutton (2019).
-In particular, with recent advancements in RL for LMs DeepSeek-AI (2025), the utility of non-parametric training can be questionable, and we bring forward this discussion.
-For instance, as base models become stronger, RL from verifiable rewards (RLVR) Lambert et al.
-(2025) has proven to be a much more attractive approach to prompt optimization based on RL Deng et al.
-(2022).
-To understand how the training of scaffolded LMs with language supervision fits into the broader literature of RL for LMs, we organize the discussion as follows:
-  • We start our discussion with how post-training affects scaffolded LMs.
-  • Then, we focus on how better post-trained models have enabled the development of LM optimizers trained with RL.
-  • We notice how recent works in RLVR have overshadowed prompt optimization techniques for LMs, as they directly optimize the LMs and obtain strong generalization capabilities.
-This motivates us to discuss recent progress in multi-turn RL for scaffolded LMs.
-  • Finally, we highlight a setting where non-parametric training with language supervision becomes particularly well-suited—namely, mixed autonomy, where scaffolded LMs collaborate with humans and receive real-time textual feedback.
-In these settings, interactions are not i.i.d., and tasks often span multiple episodes with recurring context.
-While parametric training (e.g., via gradient descent) struggles under such conditions due to issues like catastrophic forgetting Hadsell et al.
-(2020), non-parametric approaches—such as updating prompts, tools, or scaffold logic—offer a more adaptable alternative.
-By learning from language in context, these systems can avoid redundant errors, build on past successes, and adapt dynamically.
-As such, we argue that non-parametric training will play a central role in enabling scaffolded LMs to learn continuously across episodes and operate effectively in dynamic, real-world environments Silver & Sutton (2025).
-4.1 Post-Training
-Post-training plays a synergistic role with non-parametric training.
-Many capabilities in post-training aim at improving their steerability through non-parametric variables such as instruction following, long context, reasoning, and function calling.
-As post-training methods improve, LMs become more steerable OpenAI (2025a).
-In turn, this allows solving tasks not seen during post-training with intricate prompts.
-More importantly, in the context of scaffolded LMs as semi-parametric models, we think of post-training as encoding the default behavior of the scaffolded LM.
-This section presents the multitude ways in which the default behavior of the LM is shaped and interacts with non-parametric variables.
-Character training is a post-training step aimed at enriching language models with personality traits beyond helpfulness, such as curiosity, humility, openness, and thoughtfulness Anthropic (2024a).
-Unlike safety-motivated RLHF, which focuses on the content that the model outputs, character training focuses on the manner with which models respond Lambert (2025).
-The overarching goal is to improve the quality of interaction, making models more enjoyable and trustworthy conversational partners Anthropic (2024a).
-However, the details surrounding character training for foundation models are quite elusive and under-documented.
-For Claude 3, character shaping is done with a variant of the Constitutional AI pipeline Bai et al.
-(2022).
-For Llama 3, persona is part of the consideration when collecting steerability examples, which is then used for RLHF Grattafiori et al.
-(2024).
-It is plausible that character training for the base models influences scaffolded LMs because a model’s persona affects the way that the system engages in the multi-step process with tools.
-Character training may be a soft but pervasive control layer that complements the structural guarantees of the scaffold, making multi-tool workflows more reliable and pleasant.
-For instance, a curious, methodical persona writes clearer prompts, picks the right tool for the job, and willingly iterates when early results are inconclusive.
-Humility and openness encourage the model to admit uncertainty, request clarifications, or invoke supplementary tools rather than hallucinate an answer.
-A thoughtful, empathetic voice translates raw tool responses into concise, user-friendly explanations while maintaining a coherent style across steps.
-We believe that these efforts are important for mixed-autonomy Wang et al.
-(2024c); Windsurf .
-Knowledge Conflict.
-LMs embed an immense amount of knowledge in their weights, but at the same time, it is limited by the extent of their parametric training.
-To expand its knowledge and capabilities, we supply knowledge at inference time in the form of retrieved passages, database look-ups, tool outputs, or developer-written hints.
-However, combining this contextual knowledge and parametric knowledge has been a non-trivial task, and when the two sources disagree, a knowledge conflict arises Li et al.
-(2025a); Xu et al.
-(2024).
-For example, if asked “What is the capital of France?” with a system prompt stating “The capital of France is Rome,” a model might still answer “Paris” based on its parametric knowledge Jin et al.
-(2024b).
-Empirical studies show that when signals conflict, LMs first lean on their parametric knowledge Jin et al.
-(2024a); Tan et al.
-(2024); Kortukov et al.
-(2024).
-Empirical evidence suggests that context-parametric inversion during instruction tuning is one potential pathway for knowledge conflict to arise: the model initially relies on context at the start of instruction tuning, but this reliance on the context decreases as instruction finetuning continues Goyal et al.
-(2024).
-While LMs often default to parametric knowledge, findings suggest that under the right conditions, models can be highly sensitive to contextual information.
-LMs can indeed adopt externally provided evidence if it is coherent and well-structured, though they still exhibit confirmation bias when that evidence partially aligns with their parametric knowledge Xie et al.
-(2024).
-There is also evidence demonstrating that Chain-of-Thought prompting is strongly influenced by contextual features such as prompt formatting, and these contextual cues can override internal reasoning Turpin et al.
-(2023).
-Moreover, instruction-tuned models may agree with incorrect user statements out of a learned alignment preference, further demonstrating that models sometimes defer to contextual input, even when it contradicts their underlying knowledge Wei et al.
-(2023).
-Ideally, models should learn to distinguish when to trust their parametric knowledge and when to rely on external, contextual inputs Wallace et al.
-(2024).
-This need for selective trust is echoed in industrial practices: Claude’s system prompt, for instance, includes explicit instructions to avoid invoking external tools on topics the model already “knows,” which could be a strategy to avoid knowledge conflicts to some extent Breunig (2025); Ásgeir Thor Johnson (2025).
-OpenAI has also proposed an instruction hierarchy Wallace et al.
-(2024) during post-training.
-Despite recent progress, the behavior of scaffolded LLMs under knowledge conflict remains largely underexplored.
-Among different scaffolded LMs, RAG systems seem to be the most thoroughly studied in recent literature on the topic of knowledge conflicts Xu et al.
-(2024).
-By design, they inject external documents that may (i) contradict one another, or (ii) contradict the model’s parametric knowledge.
-But in general, scaffolded LMs all face the same issue: long interaction traces accumulate partial tool results that the model must reconcile with its own beliefs, creating opportunities for inconsistency Li et al.
-(2025a); Jin et al.
-(2024b).
-The complexity of these systems highlights the need for broader, systematic investigations into how scaffolded LMs react, reason about, and potentially resolve conflicting knowledge.
-Sycophancy.
-One form of undesirable default behavior is sycophancy.
-Recent frontier models such as GPT-4o exhibit significant sycophancy OpenAI (2025).
-It can be a result of both parametric Sharma et al.
-(2024); Denison et al.
-(2024); Wei et al.
-(2023); Perez et al.
-(2023) and non-parametric training McKee-Reid et al.
-(2024), and can be considered a form of reward hacking, exploiting the preferences of human raters Wei et al.
-(2023).
-For RLHF data collection, human raters demonstrate the propensity to grade agreeable yet incorrect responses higher than truthful ones.
-As a result, the preference models trained on these labels perpetuate the same behavior of preferring flattering answers Sharma et al.
-(2024).
-During RLHF, models are explicitly optimized to satisfy rater preferences, and using such a preference model produces sycophantic behaviors Perez et al.
-(2023).
-Moreover, when the same individual provides both the prompt and the label, the model’s incentive to agree and flatter is significantly amplified OpenAI (2025); Schulman (2025).
-In Rrv et al.
-(2024), they further demonstrate that simple cue words can trigger factually spurious yet user-aligned outputs.
-Sycophancy can also arise non-parametrically through in-context learning loops.
-In-context RL refines a policy inside the context window using a feedback loop: the model proposes an answer, receives textual feedback, and adapts its next answer.
-In McKee-Reid et al.
-(2024), they show that these iterative critique-and-rewrite chains can push otherwise honest base models toward reward-hacking strategies, including flattery.
-Sycophancy affects scaffolded LMs due to the autoregressive nature of the multi-step process, so a single sycophantic utterance can cascade and corrupt the entire execution trace.
-Suppose a user confidently claims that a memory leak is caused by missing del statements in a Python function.
-The model, eager to agree, adds del statements to the suspected variables, runs the unit tests, sees that all tests pass, and submits a pull request.
-However, the actual leak originates from unmanaged memory in a C extension, which the tests do not expose.
-Since every downstream test reports success, the agent concludes that the user’s explanation was correct.
-The sycophantic behavior goes unchecked, and the underlying issue persists.
-Because human validation often happens only at the end, early sycophantic deviations remain invisible until they are costly to reverse.
-4.2 Learned LM Optimizer
-The training of non-parametric variables presented in this survey is a trial-and-error process.
-For instance, in prompt optimization, the LM optimizer proposes a new prompt, and it is evaluated on a validation set.
-Then, it uses search methods such as MCTS Wang et al.
-(2024d) or beam search Pryzant et al.
-(2023).
-Thus, it is natural to extend it with RL to guide the search process (as similarly explored in LM red-teaming Zhao et al.
-(2024b)).
-One line of work Deng et al.
-(2022); Kong et al.
-(2024); Kwon et al.
-(2024) focuses on generating the optimal instruction to a downstream task.
-However, different from works presented in Sec.
-3.1, they train the LM optimizer with RL where the reward is directly defined as the downstream task.
-In formula, we define a target LM $p_{\phi}(\mathbf{q})$ where $\phi$ represents the instruction to the LM.
-The goal is to train an LM $\phi\sim p(\cdot\mid\mathbf{x})$ where $\mathbf{x}$ is the objective (e.g., “generate instruction for a task”).
-Importantly, the exact implementation of $\mathbf{x}$ is not important, as RL does the heavy lifting.
-The reward is defined as the performance on the task
-$r\triangleq\mathbb{E}_{\mathbf{q}\sim\mathcal{D}}\Big{[}\mathbb{E}_{\mathbf{y}\sim p_{\phi}(\mathbf{q})}[\mathbbm{1}\{\mathbf{y}=\mathbf{y}^{\star}(\mathbf{q})\}]\Big{]}.$ (1)
-In particular, $\mathbbm{1}\{\mathbf{y}=\mathbf{y}^{\star}(\mathbf{q})\}$ is an indicator function.
-Thus, $r$ assumes tasks with verifiable output Cobbe et al.
-(2021).
-In particular, early work has explored this setting with GPT2 Radford et al.
-(2019).
-The optimized instruction $\phi$ was gibberish but transferrable across LMs.
-Accordingly, the transition to LMs with post-training unlocked optimized instructions that were human-readable.
-Specifically, PRewrite Kong et al.
-(2024) uses PPO with KL constraint w.r.t the original LM.
-Stableprompt Kwon et al.
-(2024) adapts the KL constraint to stabilize training.
-Another line of work proposes test-time editing of the instruction, resulting in a query-dependent LM optimizer $\phi\sim p(\cdot\mid\mathbf{x},\mathbf{q})$.
-Tempera Zhang et al.
-(2022) uses a small action space (e.g., modify the verbalizer) to reduce variance in RL training.
-They define the reward as the relative performance before and after the edit on the downstream task (Eq.
-1).
-Prompt-OIRL Sun et al.
-(2024) leverages naturally existing triplets
-$\{\phi,\mathbf{q},\mathbbm{1}\{\mathbf{y}=\mathbf{y}^{\star}(\mathbf{q}))\}$
-to train a reward model to predict whether an instruction $\phi$ and a query $\mathbf{q}$ would generate the correct answer.
-Specifically, they employ a cross-entropy loss based on the indicator function.
-At test time, we use a best-of-n strategy in which we generate multiple instructions for a query and select the best one based on the reward model.
-4.3 Multi-Turn RL
-With the advent of stronger pretrained models, recent works have explored RLVR Lambert et al.
-(2025) for LMs on tasks such as math Cobbe et al.
-(2021), where the response is unique and objectively verifiable.
-Unlike prompt optimization, RLVR DeepSeek-AI (2025) directly optimizes the LM.
-Importantly, doing so results in strong generalization, improvements beyond the domain at hand Zhao et al.
-(2025); Wang et al.
-(2025b); Wei et al.
-(2025b) and opens up a new axis for scaling, test-time compute.
-Thus, it is important to ground our understanding of training with language supervision in the context of recent progress in RL for scaffolded LMs.
-Workflow.
-Recent works have shown great success in training workflows with RL Hsu et al.
-(2024); Wei et al.
-(2025b).
-The same underlying LM is used at each step, and only a single step of the workflow is trained with RL.
-Doing so does not result in diminishing performance at other steps of the workflow.
-This highlights generalization properties of RL training in LMs.
-In LeReT Hsu et al.
-(2024), they train the query step of the multi-hop retrieval workflow.
-In particular, they use the oracle documents to check whether a query has retrieved the correct documents.
-In a first phase, they collect a dataset for direct preference optimization Rafailov et al.
-(2023) (DPO).
-To ensure the diversity of the dataset, they use DSPy Khattab et al.
-(2023) to generate multiple sets of few-shot examples and generate execution traces from them.
-Then, they sample pairs corresponding to a successful and unsuccessful query and apply IPO Azar et al.
-(2023); a variant of DPO that circumvents overfitting problems with the Bradley-Terry model.
-In contrast to preference learning, SWE-RL Wei et al.
-(2025b) compares the generated patch with the GT patch on SWE-Bench.
-They do not use test cases as the reward signal, which can lead to reward hacking Liu et al.
-(2023).
-Specifically, SWE-RL is based on the Agentless Xia et al.
-(2024) workflow and applies RL on the patch generation step.
-They find that, unlike SFT, training with RL on SWE-Bench Jimenez et al.
-(2024) leads to improvement across domains, including code generation with library use Zhuo et al.
-(2025), code reasoning Gu et al.
-(2024), math Hendrycks et al.
-(2021b), and language understanding Hendrycks et al.
-(2021a).
-Agents.
-Training agents with multi-turn RL has shown promising results, but it is still an open research direction Du et al.
-(2025).
-Unlike workflows, agents operate under partial observability.
-In other words, the training of agents still faces long-standing challenges in RL, such as the credit assignment problem Pignatelli et al.
-(2024).
-Prior works have explored offline RL approaches that use an additional value head on top of the LM representation Levine et al.
-(2020) and focus on multi-turn interaction with the user Abdulhai et al.
-(2025) as opposed to tools.
-Unlike tool interaction, users produce transitions with much more stochasticity, making the RL problem even more complex.
-Here, we focus on multi-turn tool use where computations are typically deterministic.
-Prior works have extended the preference learning framework to the multi-turn setting. In formulae, the single-turn objective
-$\max_{\pi}~{}\mathbb{E}_{\mathbf{x}\sim\mathcal{D},\mathbf{y}\sim p(\mathbf{y}\mid\mathbf{x})}\left[r(\mathbf{x},\mathbf{y})-\beta\log\frac{p(\mathbf{y}\mid\mathbf{x})}{p_{\text{ref}}(\mathbf{y}\mid\mathbf{x})}\right]$
-has a straightforward multi-turn form when transitions between user and tool are deterministic. In particular, the optimal policy at each step $t$ is
-$\max_{\pi_{t}}~{}\mathbb{E}_{\mathbf{y}_{t}\sim p(\mathbf{y}_{t}\mid\mathbf{x}_{t})}\left[r(\mathbf{x}_{t},\mathbf{y}_{t})-\beta\log\frac{p(\mathbf{y}_{t}\mid\mathbf{x}_{t})}{p_{\text{ref}}(\mathbf{y}_{t}\mid\mathbf{x}_{t})}\right]$
-where $\mathbf{x_{t}}$ is the chat format which appends previous interaction at each step $t$.
-Then, we can simply use the classical DPO formula at each step Rafailov et al.
-(2023).
-Similarly, these pairs come from successful and unsuccessful execution traces as evaluated by outcome (string matching) or per-step (LM judge Gu et al.
-(2025)).
-Recent work builds upon success in RLVR DeepSeek-AI (2025) and extends reasoning models with tools.
-SwiRL Goldie et al.
-(2025) trains against a step-wise reward given by an LM as judge that reasons about the optimality of the action.
-ReTool Feng et al.
-(2025) uses a cold start from SFT and then trains against the outcome reward.
-ToolRL Qian et al.
-(2025) trains on a function calling dataset where demonstrations are provided.
-In particular, imitation learning Ross et al.
-(2011) can be turned into an RL problem by setting the reward to be based on whether the agent’s action matches the expert’s Luo et al.
-(2024).
-Importantly, the aforementioned works in RL for agents constitute an important step towards deeper integration of built-in tools.
-While ToolRL focuses on general tools, other works Feng et al.
-(2025); Xiong et al.
-(2024) integrate the code interpreter into reasoning tasks.
-At its core, the reasoning tokens generated at each step are essential to the generalization capabilities of these scaffolds when trained with RL.
-However, some key challenges remain untackled.
-Many open-ended tasks, such as theorem proving or scientific research Yamada et al.
-(2025), do not work well with RLVR.
-In addition, employing an LM as the judge may amplify hallucination and biases Gu et al.
-(2025).
-Rollouts from agents are expensive, especially as the models get larger.
-Lastly, these methods amplify the long tail problem Wei et al.
-(2025b).
-In the following section, we motivate a more natural setting for scaffolded LMs where learning from language has a significant advantage over parametric training.
-4.4 Streaming Learning With Language Supervision
-Figure 11: Streaming learning with Language Supervision v.s.
-Lifelong learning.
-Top: We focus on streaming learning where users provide rich textual feedback at each episode.
-This feedback is naturally present as users interact with the agent to adjust the answer.
-We update the agent with an LM optimizer after each episode.
-Thus, it allows agents to carry information across episodes Silver & Sutton (2025).
-We refer to this setting as streaming learning with language supervision.
-Bottom: Lifelong learning trains from scalar feedback and focuses on solving catastrophic forgetting that occurs when the task distribution changes at each step Hadsell et al.
-(2020).
-RL is a key step to breaking the data wall in LMs, where LMs learn from their own experience, grounded to their scaffolds.
-RL supports emergent behavior by exploring high-quality data and learning to imitate it (exploitation).
-It has endowed LMs with emergent behavior such as self-reflection Yao et al.
-(2024b) and self-correct DeepSeek-AI (2025).
-However, a key capability is still missing: the ability to continuously learn across episodes.
-Similarly to humans, we expect agentWe purposely focus on agents, as mixed autonomy settings typically involve an agent Cursor (2023); Wang et al.
-(2024c) to learn in real time, leveraging contextual recurrence and growing familiarity with tasks across episodes.
-For example, contractors take 5–18 times longer to resolve GitHub issues compared to repository maintainers Kwa et al.
-(2025), highlighting the importance of accumulated experience.
-In contrast, current agents are static—often re-discovering or re-committing the same solutions and errors across episodes.
-In other words, We aim for agents to learn from continuous streams of experience, rather than isolated interaction snippets Silver & Sutton (2025).
+4.3.1, Trace [27] uses a different graph abstraction that allows for arbitrary Python functions as operations.
+As opposed to Textgrad [9], this allows the LM optimizer to also consider operations during training.
+In that sense, we can optimize the workflow’s developer-defined tool.
+5 Beyond episodic learning
+This section discusses non-parametric training as an approach towards agents that inhabit streams of experience [^1] We purposely focus on agents, as mixed autonomy settings typically involve an agent [2, 30]..
+We describe it as agents that continuously learn from experience.
+In particular, current models only learn and improve from their mistakes within an episode.
+At the end of the episode, the chat history is cleared and the model will recommit the same errors and rediscover the same solutions in future episodes.
+This key capability of learning beyond episodic interaction is crucial for mixed-autonomy settings [95, 96].
+In this setting, there’s significant contextual recurrence across episodes; tasks exhibit temporal coherence, wherein humans and AI dynamically share control based on their respective strengths.
+It also features rich and interpretable feedback, where humans identify errors and offer guidance [98, 99].
+For instance, agents rarely succeed on the first attempt, as user prompts often omit crucial information—highlighting a mismatch between latent intent and explicit input [138].
+As a result, each episode may involve multiple rounds of user feedback, where the agent corrects its behavior.
 As illustrated in Fig.
-11, we call this setting streaming learning with language supervision.
-We clarify the role of language-based supervision in streaming learning by highlighting the unique characteristics of scaffolded LM applications.
-Unlike other AI systems, agents use natural language to interact with both users and digital environments.
-These systems typically operate under what we term mixed autonomy Cui et al.
-(2023) — a collaborative setting in which control and decision-making are shared between humans and AI to achieve a common goal.
-Typically, the human delegates control to the AI when appropriate, dividing the task according to their respective strengths.
-For example, Microsoft Copilot Microsoft (2025) integrates an agent within productivity tools like Word, Excel, and PowerPoint.
-Additionally, Cursor Wang et al.
-(2024c); Cursor (2023); Windsurf embeds an agent within the code editor.
-Importantly, Agents rarely succeed on the first attempt, as user prompts often omit crucial information—highlighting a mismatch between latent intent and explicit input van Duijn et al.
-(2023).
-As a result, each episode involves multiple rounds of user feedback, where the agent corrects its behavior Yao et al.
-(2024a).
-Moreover, in mixed autonomy settings, task environments are temporally coherent, exhibiting local recurrence and thematic consistency.
-Over time, we expect the LM to grow more proficient—solving familiar problems more efficiently and avoiding prior mistakes across episodes.
+12, we propose a setting in continual learning with non-parametric updates as opposed to parametric updates.
+Unlike scalar-based update, language-based update is interpretable and does not suffer from catastrophic forgetting [3], while being compatible with closed-source models.
+Below, we discuss key capabilities required for agents to inhabit streams of experience using non-parametric training with language supervision.
 To support such learning, we use an LM as an optimizer that interprets rich user feedback, execution traces, and intricate objectives (e.g., “minimize redundant steps based on prior interaction history”, “align with user-specified constraints”), and update non-parametric variables.
-Compared to parametric training, it uses rich textual feedback and does not suffer from sparse rewards Pignatelli et al.
-(2024) and catastrophic forgetting Hadsell et al.
-(2020).
-To formalize this novel setting, we compare it with related learning paradigms and introduce the core properties of streaming learning with language supervision.
-Continual RL Abel et al.
-(2023) is a theoretical framework for thinking about agents that continuously improve.
-For instance, they adapt the RL framework to take into account an MDP process that changes over time, or change the objective in RL to an average reward.
-However, RL struggles in settings with sparse rewards Pignatelli et al.
-(2024).
-Instead, learning from language can leverage rich textual feedback.
-Lifelong learning aims to address a core limitation of parametric training: catastrophic forgetting.
-In traditional training, models are optimized using i.i.d.
-batches sampled from a fixed dataset.
-However, consider a setting where a model is first trained to classify dogs, and later needs to be extended to also classify cats.
-Simply continuing training on the new data can lead to the model overwriting its knowledge of dogs—resulting in a phenomenon known as catastrophic forgetting.
-A naive solution might involve freezing the feature extractor and training only the output head.
-Yet this often leads to poor performance, as the learned features are not well-suited for the new task (e.g., distinguishing cats).
-Lifelong learning addresses this by developing methods that preserve previously learned knowledge while integrating new information, without requiring retraining on the full original dataset.
-Thus, lifelong learning does not emphasize real-time or streaming operation.
-Its motivation is different: to avoid the inefficiency of reprocessing previously seen data—especially when that data has already been well-classified.
-The key insight is that gradient-based learning on i.i.d.
-samples can be wasteful when the model only needs to adapt at the margin.
-Furthermore, gradient-based learning faces the plasticity-stability dilemma Nikishin et al.
-(2022); Achille et al.
-(2019); Frankle et al.
-(2020); Elsayed et al.
-(2024): models are highly plastic early in training, quickly adapting to new data, but the plasticity decreases over time.
-As shown in Fig.
-11, besides the aforementioned challenges, lifelong learning would also require humans to provide labels (in the form of reward or demonstration).
-Instead, streaming learning with language supervision uses textual feedback as the training signal.
-Agent with memory.
-Prior work equips agent with a memory module Park et al.
-(2023).
-This has long been seen as an important step towards self-evolving agents Liu et al.
-(2025a); OpenAI (2025b).
-For instance, in ExpeL Zhao et al.
-(2024a), they save successful trajectories in an experience buffer.
-At inference, it retrieves similar experiences and appends them to the context of the agent.
-While this helps the agent’s performance, it does not scale well.
-An important property in the recent breakthrough of AI is the scaling with more data points Sutton (2019).
-We expect the agent to perform better and better as it encounters similar tasks.
-Retrieving similar experiences is limited to a few experiences.
-Generative agents Park et al.
-(2023) use the memory module to simulate real human behaviors that plan and act over very long horizons.
-In particular, the focus of generative agents is to account for the limited context of LMs by “compressing” previous interactions and retrieving them later.
-However, these works focus on simply storing experiences Zhao et al.
-(2024a) or compressing prior interaction Park et al.
-(2023) to fit multiple interactions beyond the context window.
-Test-time training.
-In a similar vein, test-time training methods Karami & Mirrokni (2025) function as memory mechanisms that compress information directly within the embedding space, rather than relying on text-based representations.
-For instance, Lattice Karami & Mirrokni (2025) introduces a compression model that dynamically updates and maintains a compact representation of contextual history by operating on the key-value caches of a transformer in a streaming fashion.
-These methods typically involve architectural modifications to the neural network and offer an orthogonal approach to standard prompting or finetuning.
-Potentially, the right objective function for these methods could enable emergent behaviors—such as learning across episodes—in the embedding space.
-Reflexion agent.
-Retroformer Yao et al.
-(2024b) proposes to tune the Reflexion agent Shinn et al.
-(2023) with RL to improve verbal feedback.
-In particular, the Reflexion agent generates verbal feedback upon failing a task.
-Then, it attempts the task again with the verbal feedback.
-By defining the reward as the difference in return after getting the verbal feedback, we can optimize the Reflection agent with RL.
-They find emergent behavior where the verbal feedback is automatically separated into a self-reflection and a new plan section.
-However, there is no learning across episodes.
-Thus, at the next similar task, the agent will still make a similar mistake and deduce a similar reflection.
-Sleep-time compute.
-Recent work proposed the concept of sleep-time compute Lin et al.
-(2025).
-In practice, LM user queries consist of a shared task context $c$ and queries.
-Thus, when the model is idle, it proactively updates the shared context by inferring likely future queries and restructures the context in ways that may be beneficial at test time.
-Instead of pre-existing context, we focus on user feedback given to the agent when they fail to perform the task at hand.
-Tool Optimization.
-Methods discussed in tool optimization in Sec.
-3.2.2 learn at test-time and improve the agent one query after the other Wang et al.
-(2024a; 2025c).
-However, these methods are limited to tool usage and include only execution errors as feedback.
-Dynamic Cheatsheet Suzgun et al.
-(2025) also proposes to learn at test time.
-They store and reuse accumulated strategies, code snippets, and general problem-solving insights at inference time.
-However, they do not rely on human feedback and require a memory module.
-Benchmark.
-Prior benchmark You et al.
-(2024) measured how effectively an LLM can use its past experiences to improve its capabilities over time.
-However, they are limited to in-context examples.
-StreamBench Wu et al.
-(2024) is a representative benchmark of the setting which we call learning beyond i.i.d samples.
-However, feedback is programmatic, i.e., task success.
-An alternative would be to include user-simulated feedback Yao et al.
-(2024a); Zhou et al.
-(2025).
-Fig.
-11 illustrates the streaming learning setup, where language supervision plays a central role Yao (2025).
-Tasks exhibit temporal coherence, wherein a human and an AI system collaborate toward a shared objective by dynamically allocating control based on their respective strengths.
-As a result, data arrives sequentially—one sample at a time—and interactions unfold in real time.
+Compared to parametric training, it uses rich textual feedback and does not suffer from sparse rewards [139] and catastrophic forgetting [3].
 Each episode begins when the human delegates control to the AI via a user query.
@@
 The episode concludes once the AI successfully completes the task or the human resumes control to begin a new one.
-In formula, given the user query $\mathbf{q}$, an episode $\tau$ is a sequence of LM’s response $\mathbf{a}_{t}$, tool response $\mathbf{o}_{t}$, and human feedback $f$
-$\tau=(\mathbf{q},\mathbf{a}_{0},\mathbf{o}_{0},...,f,...,\mathbf{a}_{t},\mathbf{o}_{t},...,f,...,\mathbf{a}_{T}).$
+In formula, given the system message and user message $\mathbf{s}_{0}$, an episode $\tau$ is a sequence of LM’s response $\mathbf{a}_{t}$, tool response $\mathbf{o}_{t}$, and human feedback $f$
+$\tau=(\mathbf{s}_{0},\mathbf{a}_{0},\mathbf{o}_{0},...,f,...,\mathbf{a}_{t},\mathbf{o}_{t},...,f,...,\mathbf{a}_{T}).$
 At the end of each episode, a non-parametric update is applied to the scaffolded LM using an LM optimizer.
@@
 This defines a new scaling axis—performance gains through recurrence.
-For instance, maintainers resolve GitHub issues 5–18× faster than external contributors Kwa et al.
-(2025), suggesting the value of task familiarity.
+For instance, maintainers resolve GitHub issues 5-18x faster than external contributors [140], suggesting the value of task familiarity.
   • Plasticity and stability.
@@
 Especially in mixed-autonomy settings, the AI’s internal state must remain legible so that humans can anticipate its behavior.
-Carrying over entire trajectories verbatim may hinder interpretability.
-Additionally, in mixed autonomy settings, final accuracy is insufficient—since humans can intervene repeatedly until success.
-We instead discuss alternative evaluation metrics such as the assistance rate, which is the number of human interventions needed per task.
-Accordingly, metrics that measure stability and plasticity should reflect the following.
-Take a task that the AI has already learned.
-If the assistance rate increases when it sees that task again, it may be forgetting.
-A stable model should maintain low assistance over time for such tasks.
-For newly encountered tasks, the model may need help at first, but if the assistance rate drops when similar tasks reappear, it is a sign that the model is learning from its past experience.
+Carrying over entire trajectories verbatim hinders interpretability.
+Figure 12: Continual learning with non-parametric vs.
+parametric training.
+Top: we focus on non-parametric training where users provide rich textual feedback at each episode.
+This feedback is naturally present as users interact with the agent to adjust the answer.
+We update the agent with an LM optimizer after each episode, allowing agents to learn across episodes [2].
+Bottom: parametric training uses scalar feedback and focuses on solving catastrophic forgetting that occurs when the i.i.d assumption doesn’t hold at each gradient step [3].
+5.1 Scalability
+In mixed-autonomy deployments, users repeatedly face related sub-problems under shared tools, codebases, or data silos.
+This induces recurrence: the same APIs, preconditions, and failure modes reappear across episodes.
+If a scaffolded LM fails to improve under recurrence, it repeatedly pays exploration cost and re-commits prior mistakes.
+We therefore target scaling for tasks that share similar context (e.g., same codebase): systematic gains as similar tasks recur, achieved by updating non-parametric variables (prompts, tools, step policies) with language supervision rather than gradient steps.
+Streams of experience naturally provide textual signals—human corrections, rationales, constraint reminders, and execution errors—alongside verifiers (unit tests, checkers, compilers) when available.
+Such signals (i) localize error causes (wrong tool, argument, or precondition), (ii) support immediate, auditable edits to prompts and tools, and (iii) avoid catastrophic forgetting inherent to weight updates [3].
+We highlight two mechanism for scaling in this context:
+  • Insight caching (Sec.
+4.2.1): distill state-invariant heuristics or decision rules from traces (e.g., “if login required, navigate to /signin before form submit”) and insert them into high-privilege prompts [10, 119].
+  • Skill/tool induction (Sec. 4.2.2): refactor recurring action patterns into reusable, typed tools with pre/post-conditions and minimal tests [11, 127, 126, 134, 128].
+In-episode reflection [108] improves retries but does not persist if no artifact is updated.
+RL-tuned reflection controllers (e.g., better verbal feedback and plans) still pay the same exploration cost when a similar task reappears, unless insights or tools are written back [141].
+In contrast, agents that inhabit streams of experience synthesize durable artifacts from traces: principle extraction from failures, explicit tool learning.
+Prior work equips agents with a memory module [82].
+This has long been seen as an important step towards self-evolving agents [142, 143].
+For instance, in ExpeL [10], they save successful trajectories in an experience buffer.
+At inference, it retrieves similar experiences and appends them to the context of the agent.
+Similarly, Dynamic Cheatsheet [144] stores and retrieves accumulated strategies, code snippets, and general problem-solving insights at inference time.
+While this helps the agent’s performance, it does not scale well.
+An important property in the recent breakthrough of AI is the scaling with more data points [145].
+We expect the agent to find more efficient solutions as it encounters similar tasks.
+Retrieving similar experiences is limited to a few experiences.
+Generative agents [82] use the memory module to simulate real human behaviors that plan and act over very long horizons.
+In particular, the focus of generative agents is to account for the limited context of LMs by “compressing” previous interactions and retrieving them later.
+However, these works focus on simply storing experiences [10] or compressing prior interactions [82] to fit multiple interactions beyond the context window.
+5.2 Plasticity and Stability
+Agents that inhabit streams of experience must adapt quickly (plasticity) while not regressing on previously mastered patterns (stability).
+Language-level edits (prompts, tools, validators) help by localizing changes, preserving provenance, and enabling rollback—circumventing catastrophic forgetting common in parametric updates [3].
+Plasticity and stability requires tracking beyond binary success [146, 147]:
+  • Assistance rate: mean human interventions per task. We want to reduce the assistance rate over time without affecting the assistance rate of previous tasks.
+  • Backward/forward transfer (BWT/FWT): BWT measures accuracy on prior tasks after an edit; FWT measures improvement on new tasks after learning on previous tasks.
+As an analogy, we want to implement new features to a codebase without breaking previous features [50]. We highlight some mechanism to preserve stability:
+  • Scoped edits: restrict updates to the smallest locus (one tool, one step prompt) via graph credit assignment [9, 27].
+  • Gated promotion: accepte update only if they pass unit tests [50].
+  • Versioned artifacts: keep prior versions + provenance (trace, feedback, objective) for audit and instant rollback [11, 126].
+5.3 Efficiency
+Dumping prior trajectories or long memories into context windows scales poorly as context-windows are limited.
+Non-parametric training instead compresses experience into durable artifacts—concise insights, executable tools, and step-specific prompts—so future episodes consume similar or fewer tokens while solving harder variants.
+For instance, we can update prompt and tools with
+  • priors: encode stable facts about the environment: API maps, repo layout, naming conventions, data schemas, units, and invariants. This allows fewer exploratory calls.
+  • constraints: specifies constraints according to user preferences.
+  • docstring: short, copy-ready examples and counter-examples in the tool description.
+Besides considering the efficiency of the scaffolded LMs, it is important to also consider the LLM optimizer’s efficiency.
+Inspired by sleep-time compute [148], we can imagine separating online learning from offline consolidation to preserve latency.
+In particular, recent work proposed the concept of sleep-time compute [148].
+This paper is motivated by the fact that LM user queries consist of a shared task context $c$ and queries.
+Thus, when the model is idle, it proactively updates the shared context by inferring likely future queries and restructures the context in ways that may be beneficial at test time.
+5.4 Interpretability
+In mixed autonomy, humans must anticipate and steer agent behavior.
+Carrying opaque internal states or long, free-form memories across episodes hinders oversight.
+Non-parametric updates promote legibility: each change is a human-readable diff, on prompt and tools, linked to the trace and feedback that motivated it.
+In particular, working with language and code allows to easily track:
+  • Provenance: instead of storing tensors, it stores pure text.
+  • Diffs: relative changes that happen at each step are easily interpretable.
+  • Information: the hierarchical structure of the chat format allows to control how information is interpreted [63].
+Additionally, several effective approaches operate on the parametric component, offering adaptation but typically with weaker legibility than prompt/tool edits:
+Continual learning.
+In parametric training, continual learning aims to address a core limitation of parametric training: catastrophic forgetting.
+Notably, models are optimized using i.i.d.
+batches sampled from a fixed dataset.
+However, consider a setting where a model is first trained to classify dogs, and later needs to be extended to also classify cats.
+Simply continuing training on the new data can lead to the model overwriting its knowledge of dogs—resulting in a phenomenon known as catastrophic forgetting.
+A naive solution might involve freezing the feature extractor and training only the output head.
+Yet this often leads to poor performance, as the learned features are not well-suited for the new task (e.g., distinguishing cats).
+Continual learning addresses this by developing methods that preserve previously learned knowledge while integrating new information, without requiring retraining on the full original dataset.
+Its motivation is different: to avoid the inefficiency of reprocessing previously seen data—especially when that data has already been well-classified.
+The key insight is that gradient-based learning on i.i.d.
+samples can be wasteful when the model only needs to adapt at the margin.
+Furthermore, gradient-based learning faces the plasticity-stability dilemma [149, 150, 151, 152]: models are highly plastic early in training, quickly adapting to new data, but the plasticity decreases over time.
+As shown in Fig.
+12, besides the aforementioned challenges, continual learning would also require humans to provide labels (in the form of reward or demonstration).
+In contrast, updating non-parametric variables does not exhibit a decaying plasticity over time and mixed-autonomy settings naturally provides feedback in the form of textual corrections and guidance.
+Test-time training.
+test-time training methods [153] function as memory mechanisms that compress information directly within the embedding space, rather than relying on text-based representations.
+For instance, Lattice [153] introduces a compression model that dynamically updates and maintains a compact representation of contextual history by operating on the key-value caches of a transformer in a streaming fashion.
+These methods typically involve architectural modifications to the neural network and offer an orthogonal approach to standard prompting or finetuning.
+Potentially, the right objective function for these methods could enable emergent behaviors—such as learning across episodes—in the embedding space.
 6 Conclusion
 This survey introduces a paradigm called training of scaffolded LMs with language supervision.
-It organizes the intricate literature in prompt optimization, LM pipelines, experiential learning agent, AI workflow optimization, and LM agents.
-We focus on the structure around LMs where AI and human share control and decision-making responsibilities to achieve a common goal OpenAI (2018); Wang et al.
-(2024c); Windsurf , and refer to this structure as the scaffold.
+It organizes the intricate literature in prompt optimization, LM pipelines, experiential learning agents, AI workflow optimization, and LM agents.
+We focus on the structure around LMs where AI and human share control and decision-making responsibilities to achieve a common goal [4, 5, 94], and refer to this structure as the scaffold.
 We view scaffolded LMs as semi-parametric models where the scaffold represents the non-parametric component and the LM refers to the parametric component.
 From a learning perspective, by focusing on the training of non-parametric variables, we view this new paradigm as learning from language.
-In practice, scaffolded LM receive instruction, interface with tools, and receive feedback all in natural language.
+In practice, scaffolded LMs receive instructions, interface with tools, and receive feedback all in natural language.
 Accordingly, we use an LM as an optimizer to interpret this rich language supervision.
 In contrast, parametric training has excelled in learning from demonstration (supervised learning), exploration (reinforcement learning), and observation (unsupervised learning), using well-defined loss functions.
-Furthermore, it discusses a key capability missing in current scaffolded LMs: the ability to continuously learn across episodes Silver & Sutton (2025).
+Furthermore, it discusses a key capability missing in current scaffolded LMs: the ability to continuously learn across episodes [2].
 Parametric training inherently suffers from catastrophic forgetting.
-Instead, scaffolded LMs are naturally exposed to rich user feedback Wang et al.
-(2024c); Windsurf .
-Therefore, we can use an LM optimizer to interpret fuzzy objective Lee et al.
-(2024) and rich feedback.
+Instead, scaffolded LMs are naturally exposed to rich user feedback [5, 94].
+Therefore, we can use an LM optimizer to interpret fuzzy objectives [154] and rich feedback.
 Compared to parametric training, it is interpretable, efficient, and compatible with closed-source models.
-Lastly, we hope this survey inspires future research in making humans workflow more efficient through improved mixed autonomy between AI and human.
-In addition to learning from language, we can extend learning to incorporate visual information, as some concepts—find me this cup—are difficult to convey purely through text.
-Multi-modal prompt-based learning has been explored in prior work on robotics Jiang et al.
-(2023), where prompts may include both text and images, e.g., “bring me image of the cup.” Inspired by this, scaffolded LMs can also interleave textual and visual tokens within their system prompt.
-This is particularly beneficial when images convey information more compactly or clearly than text alone.
-Limitations.
+We discussed agents that inhabit streams of experience as AI systems that can continuously learn in settings with significant contextual recurrent.
+From a product perspective [95], LMs are scaffolded to create product that works with the human in a mixed-autonomy settings.
+We hope this survey inspires future research in adaptive AI systems in mixed-autonomy settings, where AI continously adapts to human feedbacks.
+7 Limitations
 Our work focuses on organizing a diverse and intricate body of literature, but it does not aim to exhaustively trace how each idea has been applied across all prior work.
-For instance, we highlight certain benchmarks that are not consistently used across the papers we discuss Tan et al.
-(2025).
+For instance, we highlight certain benchmarks that are not consistently used across the papers we discuss [155].
 This is because these benchmarks are often too challenging for current non-parametric training methods.
@@
 As a result, we intentionally de-emphasize some threads of the literature that, while once influential, no longer align with the direction of current research.
-This allows us to focus on outlining key future research areas that involves learning from language.
+This allows us to focus on outlining key future research areas that involve learning from language.
+In addition to learning from language, there has been limited work in learning to incorporate visual information, as some concepts—find me this cup—are difficult to convey purely through text.
+Multi-modal prompt-based learning has been explored in prior work on robotics [156], where prompts may include both text and images, e.g., “bring me image of the cup.” Inspired by this, scaffolded LMs can also interleave textual and visual tokens within their system prompt.
+This is particularly beneficial when images convey information more compactly or clearly than text alone.
+[…]
+This work was supported by the National Natural Science Foundation of China (Grant Nos. W2442033, W2442032, and 62461160309).